### v1

# Denoising Diffusion Probabilistic Models for Coastal Inundation Forecasting

Abstract
Coastal flooding poses significant risks to communities, necessitating fast and accurate forecasting methods to mitigate potential damage. To approach this problem, we present Diff-Flood, a probabilistic spatiotemporal forecasting method designed based on denoising diffusion models. Diff-Flood predicts inundation level at a location by taking both spatial and temporal context into account. It utilizes inundation levels at neighboring locations and digital elevation data as spatial context. Inundation history from a context time window, together with additional co-variates are used as temporal context. Convolutional neural networks and cross-attention mechanism are then employed to capture the spatiotemporal dynamics in the data. We trained and tested Diff-Flood on coastal inundation data from the Eastern Shore of Virginia, a region highly impacted by coastal flooding. Our results show that, Diff-Flood outperforms existing forecasting methods in terms of prediction performance ($6\%\text{ to }64\%$ improvement in terms of two performance metrics) and scalability.
1 Introduction
The projection in sea level rise have highlighted the increasing vulnerability of many coastal regions to inundation (Wuebbles et al., 2017). Among these regions, the East Coast in US is a region of particular policy concern due to its geography and low-lying topography (Ezer and Atkinson, 2014). This region, in recent periods, has been experiencing frequent inundation events  (Ezer, 2020), causing disruptions in daily life (e.g. traffic, economic) (Jacobs et al., 2018). In the future, this situation will only keep declining due to the combined effect of climate change, sea level rise, and urbanization (Swain et al., 2020). Therefore, forecasting coastal inundation is of utmost importance from a policy standpoint for undertaking risk prevention measurement or safe evacuation of the residents in danger  (Yang et al., 2019).
Figure 1: Flooding on Quinby Bridge Road, Accomack County, VA on April 3, 2024 (Data source: Tidewatch Maps (Tidewatch, 2025).)
The problem of coastal inundation forecasting came to our attention during conversations with staff at the City of Virginia Beach. The region faces challenges with coastal flooding which causes slow-down and/or road closures; disrupting the day-to-day activities of the inhabitants. A fast and high-resolution flood forecasting model would enable city officials to identify localized inundation risks, resulting in more effective preparedness and response. In this paper, we approach the problem of designing such a model. It is a joint-work of a multidisciplinary team, including computer and data scientists, academic experts with expertise in hydro-dynamic models, members of established coastal resources management centers, and officials at the City of Virginia Beach.
The traditional method for predicting inundation involves hydro-dynamical, physics-based models (PBM). Commercially available (Syme, 2001; Deltares, 2014) (e.g., SOBEK, TUFLOW) or open-sourced (Zhang et al., 2016) (e.g, SCHISM) hydro-dynamical models can simulate inundation accurately at a high resolution. However, the computation time associated with these PBMs makes them challenging for real-time forecasting (Guo and others, 2021). Although researchers have investigated Deep-Learning (DL) based time-series prediction methods as surrogates to these PBMs (Roy et al., 2023), these models are not suitable for accommodating spatial contexts. Existing spatiotemporal forecasting methods, on the other hand, struggle to scale in high-resolution forecasting tasks.
Motivated by the above observations, we explore the Diffusion Model (Sohl-Dickstein et al., 2015) in the context of high-resolution coastal inundation forecasting task. The primary reason for investigating diffusion models is its ability to utilize conditioning contexts of various dimensions (Rombach et al., 2022; Zhang and others, 2023) and its potential to model physical properties (Ahmad and others, 2024; Amram and Pedro, 2023). Originally developed for image generation, diffusion models have been adapted for time-series forecasting, achieving state-of-the-art performance, as demonstrated by Rasul et al. (2021). However, their approach does not incorporate spatial context. More recently,  (Wen et al., 2023) proposed a probabilistic spatiotemporal graph forecasting method based on denoising diffusion models. While promising, such methods face scalability challenges when applied to large-scale/high-resolution inundation forecasting tasks. Using the coastal inundation of the Eastern Shore as case study, we address these challenges by making the following contributions in this paper.
  • We formulate the problem of coastal inundation forecasting under the spatiotemporal forecasting framework and introduce Diff-Flood as a novel method for approaching the problem. It can utilize both spatial and temporal context to generate spatiotemporal forecasts at fine spatial resolution.
  • Extensive evaluation of our method against several baseline methods on inundation data of the Virginia Eastern Shore underscores the superiority of our proposed method and its aspect of being used as a decision-making tool for policymakers. Specifically, we observed $6\%$ to $64\%$ improvement in prediction performance, in terms of two performance metrics (Normalized Average Continuous Ranked Probability Score, and Normalized Root Mean Square Error). An ablation study of our model is performed to highlight the importance of its components.
  • Furthermore, we demonstrate the scalability of our model compared to other baselines in large settings, underscoring its utility in real-time forecasting in large-scale scenarios.
It is worth highlighting that our model can be used to answer inundation scenario-based queries relevant to policymakers. For example, queries of the form: “What is the probability that a particular region will get inundated by 1 feet within the next 12 hours?”, can be answered by using our model through probabilistic forecasting, see section 7.
2 Related Work
We present an abridged version of the most relevant literature in this section. The supplementary material of this paper covers the literature more extensively.
Hydrology Models: Traditionally, various Physics Based hydrodynamic Models (PBM) have been employed to simulate inundation. SOBEK (Deltares, 2014) is a 2D model that solves a Saint-Venant flow equation and shallow water equation to determine the water level of rectangular grids. TELEMAC-2D (Vu et al., 2015) uses a similar method but is more suitable for coastal areas. Polymorphic models like SCHISM (Zhang et al., 2016) can capture water flow at very high resolution. However, the computation time associated with solving these complex equations at high spatial resolution makes these PBMs infeasible for real-time forecasting.
Deep Learning: DL methods are a popular choice in surrogating hydrodynamic models. Regression-based methods (Zahura and Goodall, 2022), recurrent neural networks (Roy et al., 2023) and hybrid ML-based methods (Zanchetta and Coulibaly, 2022) are some of the methods that have been attempted as possible surrogates. However, these models cannot inherently take spatial context into account. The DCRNN method proposed by (Li et al., 2017) for traffic forecasting is the state-of-the-art DL method for spatiotemporal point forecasting, where bidirectional random walk is used to capture spatial dependency and Seq2Seq architecture is used to capture temporal dependency.
Diffusion Model The field of generative modeling has been dominated by the diffusion model for quite some time. Although it had primarily been studied for image synthesis (Ho et al., 2020; Dhariwal and Nichol, 2021; Austin et al., 2021), its fascinating ability to use flexible conditioning to guide the generation process has been utilized in other domains like video generation (Yang et al., 2023b), prompt-based image generation (Ramesh et al., 2022), text-to-speech generation (Yang et al., 2023a). TimeGrad (Rasul et al., 2021) was the pioneering method that approached the time-series forecasting problem by utilizing a diffusion model. Other models like CSDI (Tashiro et al., 2021), DPSD-CSPD (Biloš et al., 2023), TDSTF (Chang et al., 2024) were later introduced. These models try to predict the noise added during the forward process of the diffusion model ($\epsilon$-parameterization). An alternative parameterization is to predict the ground-truth data ($x$-parameterization). Different from models mentioned above, $D^{3}$ (Li et al., 2022) uses coupled diffusion process and bidirectional variational autoencoder to better forecast in longer horizon and improve interpretability. The State-of-the-Art Diffusion-based model DiffSTG (Wen et al., 2023) devised a modified architecture for the reverse process of diffusion model combining UNet and GNN to capture temporal and spatial dependencies, respectively. However, both these methods scale poorly with the growth in number of variables.
3 Problem Formulation
Formally, we define the inundation forecasting problem as follows.
Problem 1.
[…]
Problem 2.
Multi-patch Probabilistic Inundation Forecasting Let $\mathcal{P}=\{P_{1},P_{2},...,P_{K}\}$ denote a set of two-dimensional patches of size ${D\times D}$. Let ${}_{k}\mathbf{x}_{t}^{0}\in\mathbb{R}^{D\times D}$ denote the inundation matrix of patch $P_{k}$ at timestep $t\in\{1,\ldots,T\}$. Let ${}_{k}\mathbf{s}\in\mathbb{R}^{D\times D}$ and ${}_{k}\mathbf{z}_{t}\in\mathbb{R}^{l},t\in\{1,\ldots,T\}$ denote the elevation matrix and time series of co-variates associated with patch $P_{k}$, respectively. Let $t_{0}:t_{c+l}$ denote the increasing sequence of timesteps ${t_{0},t_{1},...,t_{c+l}}$, where $c,l>0$. We aim to learn the conditional distribution function $q({}_{k}\mathbf{x}_{t_{c}:t_{c+l}}^{0}|{}_{k}\mathbf{x}_{t_{0}:t_{c-1}}^{0},{}_{k}\mathbf{z}_{t},{}_{k}\mathbf{s})$.
Note that, both problems are spatio-temporal forecasting problems, since we aim to forecast the inundation level on all cells in a patch together. They are also probabilistic forecasting problems, as we want to learn the distribution of inundation values instead of making a single prediction. Finally, Problem 2 represents a generalized version of Problem 1. Here, we aim to learn a single model that can be utilized to forecast inundation levels on multiple patches/areas.
The following section presents the basics of a denoising diffusion model used by our method Diff-Flood for learning our target distributions. Subsequently, we present details on how we have designed Diff-Flood to capture spatial and temporal dynamics of our inundation values.
4 Denoising Diffusion Model
In this section, we provide a brief overview of denoising diffusion models. A more comprehensive view is provided in the supplementary material.
[…]
$q(\mathbf{x}^{1:N}|\mathbf{x}^{0})=\prod_{n=1}^{N}q(\mathbf{x}^{n}|\mathbf{x}^{n-1})$
is fixed to a Markov chain (often referred to as the forward process), and $(iii)$ the structure of the latent encoder at each hierarchical step is pre-defined as a linear Gaussian model as follows:
$q(\mathbf{x}^{n}|\mathbf{x}^{n-1})=\mathcal{N}(\mathbf{x}^{n};\sqrt{1-\beta_{n}}\mathbf{x}^{n-1},\beta_{n}\mathbf{I}).$
Additionally, the Gaussian parameters ($\beta_{1},\beta_{2},...,\beta_{n})$ are chosen in such a way that we have: $q(\mathbf{x}^{N})=\mathcal{N}(\mathbf{x}^{N};{\mathbf{0}},\mathbf{I})$.
The joint distribution $p_{\theta}(\mathbf{x}^{0:N})$ is called the reverse process. It is defined as a Markov chain with learned Gaussian transitions beginning at $p(\mathbf{x}^{N})$:
$p_{\theta}(\mathbf{x}^{0:N})=p(\mathbf{x}^{N})\prod_{n=1}^{N}p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})$
Each subsequent transition is parameterized as follows:
$p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})=\mathcal{N}(\mathbf{x}^{n-1};\mu_{\theta}(\mathbf{x}^{n},n),\Sigma_{\theta}(\mathbf{x}^{n},n)\mathbf{I})$ (1)
Both $\mu_{\theta}:\mathbb{R}^{D}\times\mathbb{N}\rightarrow\mathbb{R}^{D}$ and $\Sigma_{\theta}:\mathbb{R}^{D}\times\mathbb{N}\rightarrow\mathbb{R}^{+}$ take two inputs: $\mathbf{x}^{n}\in\mathbb{R}^{D}$ and the noise step $n\in\mathbb{N}$. The parameters $\theta$ are learned by fitting the model to the data distribution $q_{\chi}(\mathbf{x}^{0})$. This is done by minimizing the negative log-likelihood via a variational bound:
[…]
$=\log\mathbb{E}_{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}\frac{p_{\theta}(\mathbf{x}^{0:N})}{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}$
$\geq\mathbb{E}_{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}\log\frac{p_{\theta}(\mathbf{x}^{0:N})}{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}$ (2)
$\implies-\log p_{\theta}(\mathbf{x}^{0})$ $\leq\mathbb{E}_{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}\log\frac{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}{p_{\theta}(\mathbf{x}^{0:N})}$ (3)
Here, we get (2) by applying Jensen’s inequality. We indirectly minimize the negative log-likelihood of the data, i.e., $-\log p_{\theta}(\mathbf{x}^{0})$, by minimizing the right hand side of (3).
Ho et al. (2020) showed that the forward process has the following property: we can sample $\mathbf{x}^{n}$ using $\mathbf{x}^{0}$ at any noise step $n$ in closed form . Let, $\alpha_{n}=1-\beta_{n}$, and $\bar{\alpha}_{n}=\prod_{i=1}^{n}\alpha_{i}$. Then, we have:
$q(\mathbf{x}^{n}|\mathbf{x}^{0})=\mathcal{N}\left(\mathbf{x}^{n};\sqrt{\bar{\alpha}_{n}}\mathbf{x}^{0},(1-\bar{\alpha}_{n})\mathbf{I}\right).$ (4)
Based on this property, Ho et al. (2020) showed that one possible parameterization of (1) is as follows:
$\mu_{\theta}(\mathbf{x}^{n},n)$ $=\frac{\sqrt{\alpha_{n}}(1-\bar{\alpha}_{n-1})}{1-\bar{\alpha_{n}}}\mathbf{x}^{n}+\frac{\sqrt{\bar{\alpha}_{n-1}}\ \beta_{n}}{1-\bar{\alpha_{n}}}\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)$ (5)
$\Sigma_{\theta}(\mathbf{x}^{n},n)$ $=\frac{1-\bar{\alpha}_{n-1}}{1-\bar{\alpha_{n}}}\beta_{n}$ (6)
Here, $\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)$ is a neural network that predicts $\mathbf{x}^{0}$ from noisy vector $\mathbf{x}^{n}$ and noise step $n$. Ho et al. (2020) then showed that, the right hand side of (3) can be minimized by minimizing:
$\mathbb{E}_{n,q(\mathbf{x}^{n}|\mathbf{x}^{0})}\ \frac{1}{2\tilde{\beta}_{n}}\frac{\bar{\alpha}_{n-1}\ \beta_{n}^{2}}{(1-\bar{\alpha}_{n})^{2}}||\mathbf{x}^{0}-\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)||_{2}^{2}$ (7)
Therefore, optimizing a diffusion model can be done by training a neural network to predict the original ground truth vector from an arbitrarily noisy version of it. Once trained, to sample from the reverse process $\mathbf{x}^{n-1}\sim p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})$, we first sample $\mathbf{x}^{N}$ from $\mathcal{N}({\mathbf{0}},\mathbf{I})$. Then, we compute:
$\mathbf{x}^{n-1}=\mu_{\theta}(\mathbf{x}^{n},n)+\sqrt{\Sigma_{\theta}}\ \mathbf{v}$ (8)
Here, $\mathbf{v}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$ for $n\in[2,N]$, and $\mathbf{v}={\mathbf{0}}$ when $n=1$. $\mu_{\theta}(\mathbf{x}^{n},n)$ and $\Sigma_{\theta}$ are computed using (5) and (6), respectively.
5 Diff-Flood Method
Diff-Flood is designed for multi-patch probabilistic inundation forecasting, where the inundation values are correlated both spatially and temporally. Following our notations in Problem 1 and 2, let ${}_{k}\mathbf{x}_{t}^{0}\in\mathbb{R}^{D\times D}$ denote the inundation matrix of the patch $P_{k}$ at time step $t$. Towards our goal of learning the conditional distribution $q_{\chi}({}_{k}\mathbf{x}_{t_{c}:t_{c+l}}^{0}|{}_{k}\mathbf{x}_{t_{0}:t_{c-1}}^{0},{}_{k}\mathbf{z}_{t_{0}:t_{c-1}},{}_{k}\mathbf{s})$, where ${}_{k}\mathbf{s}\in\mathbb{R}^{D\times D}$ is the elevation matrix of patch $P_{k}$, and ${}_{k}\mathbf{z}_{t}$ denotes a time-series of co-variates associated with $P_{k}$, we assume:
$q_{\chi}({}_{k}\mathbf{x}_{t_{c}:t_{c+l}}^{0}|{}_{k}\mathbf{x}_{t_{0}:t_{c-1}}^{0},{}_{k}\mathbf{z}_{t_{0}:t_{c-1}},{}_{k}\mathbf{s})=$
$\quad\quad\quad\textstyle\prod_{t=t_{c}}^{t_{c+l}}q_{\chi}({}_{k}\mathbf{x}_{t}^{0}|{}_{k}\mathbf{x}_{t_{0}:t-1}^{0},{}_{k}\mathbf{z}_{t_{0}:t-1},{}_{k}\mathbf{s})$ (9)
In Diff-Flood, we use a denoising diffusion model (as described in Section B) to learn the conditional distributions on the right hand side of (9). We refer to ${}_{k}\mathbf{x}_{t_{0}:t-1}^{0}$, ${}_{k}\mathbf{z}_{t_{0}:t-1}$, and ${}_{k}\mathbf{s}$ as context data.
Figure 2: Architecture of proposed Diff-Flood Model
To capture spatial correlation, we first augment the inundation data in the context (i.e. ${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$) by concatenating the elevation matrix ${}_{k}\mathbf{s}$ to it as an additional channel (Equation 10). The goal is to capture the correlation of elevation and inundation values. We then use convolution blocks extract spatial features ${}_{k}\mathbf{f}_{t_{0}:t-1}$ (Equation 11).
${}_{k}\mathbf{y}^{0}_{t_{0}:t-1}={}_{k}\mathbf{x}^{0}_{t_{0}:t-1}\oplus{}_{k}\mathbf{s}$ (10)
${}_{k}\mathbf{f}_{t_{0}:t-1}=\textsc{conv}{}_{\theta_{1}}({}_{k}\mathbf{y}^{0}_{t_{0}:t-1})$ (11)
In addition, since coastal inundation is expected to be correlated with tide cycles, we use the hour of day, and day of month as temporal features $\mathbf{z}_{t}$. We use sinusoidal encoding to capture the cyclic nature of these temporal features.
${}_{k}\mathbf{g}^{0}_{t_{0}:t-1}=\textsc{sinusoid}{}({}_{k}\mathbf{z}_{t_{0}:t-1})$ (12)
We concatenate the spatial and temporal features, and then pass it through a linear fully connected layer to get an embedding of the context.
${}_{k}\mathbf{h}_{t_{0}:t-1}=\textsc{linear}{}_{\theta_{2}}({}_{k}\mathbf{f}_{t_{0}:t-1}\oplus{}_{k}\mathbf{g}_{t_{0}:t-1})$ (13)
An overview of the architecture of Diff-Flood is provided in Figure 2. We approximate (9) by the model in (14). Here, $\theta$ comprises of the trainable parameters of the convolution blocks ($\theta_{1}$), fully connected layer ($\theta_{2}$), and the denoising diffusion model ($\theta_{3}$).
$\textstyle\prod_{t=t_{c}}^{t_{c+l}}p_{\theta}({}_{k}\mathbf{x}_{t}^{0}|{}_{k}\mathbf{h}_{t_{0}:t-1})$ (14)
As described in Section B, within the denoising diffusion model, we need to train a neural network $\hat{\mathbf{x}}_{\theta_{3}}({}_{k}\mathbf{x}^{n}_{t},n,{}_{k}\mathbf{h}_{t_{0}:t-1})$ that predicts ${}_{k}\mathbf{x}^{0}_{t}$ from noisy vector ${}_{k}\mathbf{x}^{n}_{t}$, noise step $n$, and context embedding ${}_{k}\mathbf{h}_{t_{0}:t-1}$. We use a conditional UNet for this purpose. Within the UNet architecture, we employ a cross-attention conditioning mechanism (Rombach et al., 2022) to map the context embedding to the intermediate layers of the UNet. It allows us to predict ${}_{k}\mathbf{x}^{0}_{t}$ while paying attention to different parts of the context embedding.
5.1 Training
Input: Data: ${}_{k}\mathbf{x}^{0}_{t}\sim q({}_{k}\mathbf{x}^{0}_{t})$, Context: ${}_{k}{\mathbf{x}^{0}_{t_{0}:t-1}}$, ${}_{k}\mathbf{z}_{t_{0}:t-1}$, ${}_{k}\mathbf{s}$.
1 while not converged do
       // Forward diffusion
2       Initialize $n\sim\text{Uniform}(1,N)$.
3       ${}_{k}\mathbf{x}^{n}_{t}\leftarrow\mathcal{N}\left(\sqrt{\bar{\alpha}_{n}}\ {}_{k}\mathbf{x}^{0}_{t},(1-\bar{\alpha}_{n})\mathbf{I}\right)$.
       // Compute context embedding
4       ${}_{k}\mathbf{f}^{0}_{t_{0}:t-1}\leftarrow\textsc{conv}{}_{\theta_{1}}({}_{k}\mathbf{x}^{0}_{t_{0}:t-1}\oplus{}_{k}\mathbf{s})$
5       ${}_{k}\mathbf{g}^{0}_{t_{0}:t-1}\leftarrow\textsc{sinusoid}{}({}_{k}\mathbf{z}_{t_{0}:t-1})$ ${}_{k}\mathbf{h}_{t_{0}:t-1}\leftarrow\textsc{linear}{}_{\theta_{2}}({}_{k}\mathbf{f}_{t_{0}:t-1}\oplus{}_{k}\mathbf{g}_{t_{0}:t-1})$
       // prediction
6       ${}_{k}\hat{\mathbf{x}}^{0}_{t}\leftarrow\hat{\mathbf{x}}_{\theta_{3}}({}_{k}\mathbf{x}_{t}^{n},n,{}_{k}\mathbf{h}_{t_{0}:t-1})$
7       Take gradient step on $\nabla_{\theta}||{}_{k}\mathbf{x}^{0}_{t}-{}_{k}\hat{\mathbf{x}}^{0}_{t}||^{2}_{2}$
Algorithm 1 Training for patch $P_{k}$, prediction time step $t$
To train Diff-Flood, we sample context windows (${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$) and the immediate next timestep (${}_{k}\mathbf{x}^{0}_{t}$) from the training time series data of each patch. We then optimize the parameters $\theta$ by minimizing the negative log-likelihood of the data ($-\log p_{\theta}({}_{k}\mathbf{x}^{0}_{t}|{}_{k}\mathbf{h}_{t_{0}:t-1}$). Through a similar derivation as shown in the Section B, the objective to minimize becomes:
$\mathbb{E}_{{}_{k}\mathbf{x}^{0}_{t},n}||{}_{k}\mathbf{x}^{0}_{t}-\hat{\mathbf{x}}_{\theta}({}_{k}\mathbf{x}_{t}^{n},n,{}_{k}\mathbf{h}_{t_{0}:t-1}||^{2}_{2}$ (15)
Here, the neural network $\hat{x}_{\theta}$ is now also conditioned on the context embedding ${}_{k}\mathbf{h}_{t_{0}:t-1}$.
Input: Noise: ${}_{k}\mathbf{x}^{N}_{t}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$, Context: ${}_{k}{\mathbf{x}^{0}_{t_{0}:t-1}}$, ${}_{k}\mathbf{z}_{t_{0}:t-1}$, ${}_{k}\mathbf{s}$.
1 for $n\in{N,...,1}$ do
2       if  $n>1$ then
3             $\mathbf{v}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$
4      else
5             $\mathbf{v}\leftarrow{\mathbf{0}}$
6      Compute context embedding ${}_{k}\mathbf{h}_{t_{0}:t-1}$.
7       ${}_{k}\hat{\mathbf{x}}^{0}_{t}\leftarrow\hat{\mathbf{x}}_{\theta}({}_{k}\mathbf{x}_{t}^{n},n,{}_{k}\mathbf{h}_{t_{0}:t-1})$
8       ${}_{k}\mathbf{x}^{n-1}_{t}\leftarrow\frac{\sqrt{\alpha_{n}}(1-\bar{\alpha}_{n-1})}{1-\bar{\alpha_{n}}}{}_{k}\mathbf{x}^{n}_{t}+\frac{\sqrt{\bar{\alpha}_{n-1}}\beta_{n}}{1-\bar{\alpha_{n}}}{}_{k}\hat{\mathbf{x}}^{0}_{t}$
9       $\hskip 160.00024pt\ignorespaces+\sqrt{\Sigma_{\theta}}\mathbf{v}$
return ${}_{k}\mathbf{x}^{0}_{t}$
Algorithm 2 Sampling ${}_{k}\mathbf{x}^{0}_{t}$
5.2 Inference
During inference, given a context ${}_{k}\mathbf{x}^{0}_{t_{0}:t-1},{}_{k}\mathbf{z}_{t_{0}:t-1},{}_{k}\mathbf{s}$, we first compute the context embedding ${}_{k}\mathbf{h}_{t_{0}:t-1}$. We then apply Algorithm 2 to obtain a sample ${}_{k}\mathbf{x}^{0}_{t}$ of the next time step. Based on this sample, we compute the new context embedding ${}_{k}\mathbf{h}_{t_{0}:t}$ and repeat until the desired number of prediction time steps have been reached. We repeat the process multiple times to get multiple predictions.
6 Experiment
We conduct our experiments in two parts. First, we evaluate Diff-Flood model on the TideWatch dataset (see Section 6.1) against five competitive baseline models, taking forecasting inundation across the Virginian Eastern Shore region as our case study. Second, we conduct an ablation study of our model, underscoring the importance of the components of our proposed architecture.
6.1 Dataset
To train Diff-Flood, we downloaded inundation data of the Eastern Shore, Virginia from Tidewatch Maps (Loftis et al., 2019; Loftis, 2022); a street-level inundation mapping tool developed by the Virginia Institute of Marine Science (VIMS). It utilizes the open-source hydro-dynamic model, SCHISM (Zhang et al., 2016) as the engine, and provides 36 hour inundation forecast maps with a 12 hour update frequency. We used this dataset due to its high spatial resolution, coverage of areas affected by coastal inundation, and lack of other publicly available inundation datasets.
We downloaded 89 days (March 28, 2024 to June 24, 2024) of hourly inundation data of the area. The data comes in a raster geo-database format with a pixel resolution of $10m\times 10m$ and provides inundation level at each pixel in feet. We resampled the data to a pixel resolution of $30m\times 30m$. We then consider different square-shaped patches of size: $16\times 16$, $64\times 64$, $80\times 80$, and $96\times 96$. Each patch $P_{k}$ has a time series of inundation levels denoted by ${}_{k}\mathbf{x}^{0}_{t}\in\mathbb{R}^{D\times D}$, where $D$ is the length of one side of the patch. We aim to train Diff-Flood to forecast inundation levels on multiple patches.
We have used 81 days (1944 timesteps, each of length 1-hour) of data for training, 1 day (24 timesteps) of data for validation, and 7 days (168 timesteps) of data for testing. Note that, when we utilize $K$ number of patches, the total number of both training and testing timesteps are multiplied by $K$. In our forecasting task, we use context window of $12$ hours, and then predict the inundation levels in the subsequent $12$ hours (prediction window). Each sample datapoint therefore contains 24 consecutive timesteps, first $12$ hours for context and the last $12$ hours for prediction. We normalize each datapoint separately as follows: first we calculate the mean and standard deviation of the inundation values in the context window, then we standardize the inundation values in both the context and prediction window using the calculated mean and standard deviation. In addition to the inundation data, we used digital elevation data (OpenTopography, 2021) for the same area. We resampled the elevation data to the same resolution of $30m\times 30m$ and then aligned it with our inundation data.
6.2 Evaluation Metric
We use two metrics to evaluate forecast quality with respect to the observations. The first metric is Normalized Root Mean Squared Error (NRMSE). Let, $\mathbf{x}$ be the set of all observations over all cells for all timesteps. $x_{ij,t}\in\mathbf{x}$ be the observed value of cell $(i,j)$ at time $t$ and let, $\tilde{x}_{ij,t}$ denote the corresponding average forecast value. Then, NRMSE is defined as follows:
$\text{NRMSE}=\frac{1}{\max_{ij,t}(\mathbf{x})-\min_{ij,t}(\mathbf{x})}\sqrt{\frac{\sum_{ij,t}{(x_{ij,t}-\tilde{x}_{ij,t})^{2}}}{\sum_{ij,t}1}}$ (16)
In addition to NRMSE, we use an extended version of the Continuous Ranked Probability Score (CRPS) Winkler et al. (1996) to quantify forecast uncertainty. The CRPS metric is calculated for each cell and each timestep individually. Let’s assume that we have $M$ forecasts ${\hat{x}_{ij,t}^{1},\hat{x}_{ij,t}^{2},...,\hat{x}_{ij,t}^{M}}$ for observation $x_{ij,t}$. First, the empirical cumulative distribution function (CDF) at a point $z$, $\hat{F}_{ij,t}^{M}(z)$ from these forecast values is defined as
$\hat{F}_{ij,t}^{M}(z)=\frac{1}{M}\sum_{m=1}^{N}{\mathbb{I}(\hat{x}_{ij,t}^{n}\leq z)}$ (17)
where $\mathbb{I}(y)$ is the indicator function that yields $1$ if condition $y$ holds, 00 otherwise. From this, CRPS between the empirical CDF $\hat{F}_{ij,t}^{M}$ and the observation $x_{ij,t}$ is calculated as
[…]
Since CRPS is defined for individual timestep and individual cells, we extend this metric to Normalized Average CRPS (NACRPS), defined as
$\text{NACRPS}(\hat{F}^{M},\mathbf{x})=\frac{\sum_{ij,t}{\text{CRPS}(\hat{F}_{ij,t}^{M},x_{ij,t})}}{\sum_{ij,t}{|x_{ij,t}|}}$ (19)
Here, $|x_{ij,t}|$ denotes the absolute value of $x_{ij,t}$.
6.3 Baseline Comparison
Patch
$D\times D\times K$
Metric DiffSTG DCRNN DeepVAR LSTNet TimeGrad Diff-Flood
$16\times 16\times 2$ NACRPS 1.2523$\pm$0.0445 N/A 1.4782$\pm$0.0332 N/A 1.2319$\pm$0.0672 0.9603$\pm$0.2119 ($\uparrow$22.05%)
NRMSE 0.1278$\pm$0.0022 0.1302$\pm$0.0125 0.1742$\pm$0.0034 0.1439$\pm$0.0003 0.1601$\pm$0.0071 0.1162$\pm$0.0189 ($\uparrow$9.02%)
$16\times 16\times 5$ NACRPS 1.3166$\pm$0.0164 N/A 1.4151$\pm$0.0293 N/A 1.2558$\pm$0.0707 0.834$\pm$0.1781 ($\uparrow$33.59%)
NRMSE 0.1373$\pm$0.0015 0.1166$\pm$0.0096 0.1734$\pm$0.0023 0.1501$\pm$0.0001 0.17$\pm$0.0075 0.1095$\pm$0.0165 ($\uparrow$6.13%)
$64\times 64\times 10$ NACRPS Failed N/A 1.0228$\pm$0.0895 N/A 0.5522$\pm$0.0912 0.2007$\pm$0.0201 ($\uparrow$63.66%)
NRMSE 0.0801$\pm$0.0065 0.3437$\pm$0.035 0.1765$\pm$0.0002 0.1818$\pm$0.0182 0.0639$\pm$0.0079 ($\uparrow$20.21%)
$64\times 64\times 20$ NACRPS Failed Failed 1.1451$\pm$0.0792 N/A 0.545$\pm$0.0181 0.227$\pm$0.0249 ($\uparrow$58.34%)
NRMSE 0.3967$\pm$0.0419 0.1941$\pm$0.0006 0.1908$\pm$0.0029 0.0794$\pm$0.008 ($\uparrow$58.39%)
$80\times 80\times 20$ NACRPS Failed Failed 1.1494$\pm$0.1409 N/A 0.5251$\pm$0.0105 0.2091$\pm$0.0216 ($\uparrow$60.18%)
NRMSE 0.2761$\pm$0.0399 0.131$\pm$0.0003 0.1251$\pm$0.0021 0.0508$\pm$0.0034 ($\uparrow$59.39%)
$96\times 96\times 20$ NACRPS Failed Failed 1.3578$\pm$0.0372 N/A 0.5788$\pm$0.016 0.2519$\pm$0.0241 ($\uparrow$56.48%)
NRMSE 0.1661$\pm$0.0057 0.0748$\pm$0.0001 0.0718$\pm$0.001 0.0334$\pm$0.0044 ($\uparrow$53.44%)
Table 1: Performance Comparison of Diff-Flood with baseline models. All metrics are reported up to one standard deviation after conducting five experiments with different random seeds. For each configuration, the best performance is bold-highlighted and the second-best performance is underlined. For Diff-Flood, percentage improvement in performance with its closest baseline competitor is shown. State-of-the-art spatiotemporal forecasting methods (e.g., DiffSTG, DCRNN) could not handle training for large patch configurations, these are marked as ‘Failed’. Since, DCRNN and LSTNet are point forecasting methods, their NACRPS metric are marked as ‘N/A’.
In this section, we describe the baseline models compared with our model. All experiments were conducted on an HPC cluster. To achieve scalability, experiments with different configurations were conducted with the SLURM scheduler. Each job used up to $256$GB CPU RAM, $8$ CPU cores and could request for one GPU with maximum $40$ GB GPU RAM.
  1. 1.
    DiffSTG (Wen et al., 2023): The State-of-the-art (SOTA) diffusion-based model for spatiotemporal forecasting.
  2. 2.
    TimeGrad (Rasul et al., 2021) The SOTA diffusion-based model designed for Multivariate time-series forecasting.
  3. 3.
    DeepVAR (Salinas et al., 2019): Combines an RNN-based architecture with a Gaussian copula process output model with a low-rank covariance structure.
  4. 4.
    LSTNet (Lai et al., 2018): The Long-and Short-term Time-series network (LSTNet) captures local dependency pattern among variables and long-term temporal correlation among time-series trends using CNN and RNN.
  5. 5.
    DCRNN (Li et al., 2017): A spatiotemporal forecasting framework that combines Diffusion Convolution and Sequence-to-sequence architecture to capture temporal and spatial dependencies.
Each model was trained on six different patch settings as shown in Table 1. All hyper-parameters of our model are provided in the supplementary material. It is to be noted that the number of variables increases quadratically with patch size and linearly with the number of patches. Since some models do not scale well [^1] It is worth mentioning we also attempted to use $D^{3}$ (Li et al., 2022) as a baseline. However, we it scaled poorly; causing memory error even for our lowest patch configuration. with increments in the number of variables, we limit $D=96$ for this set of experiments. Across all experiments, we set the context and prediction length to 12. Eight scenarios were sampled to understand the quality of probabilistic forecasting. Table 1 summarizes the performance of our model with respect to the baseline models. Below, we outline some key observations from the results.
Figure 3: Scaling in Model Parameters (Log scale in Y axis) with increasing patch configuration. It has to be noted that model parameter is not affected by $K$. DiffSTG and DCRNN are not shown here since they fail to execute for large patch setting.
First, in terms of NACRPS, our model consistently outperforms the baselines, underscoring its superiority in probabilistic forecasting. For small patch size ($D=16$), Diff-Flood has a performance gain of between $22.05\%$ and $33.59\%$, compared to its closest baseline. For larger patch sizes ($D\geq 64$), the improvement in NACRPS is close to $60\%$, underscoring the performance scalability of our model. Second, Diff-Flood outperforms other baselines in terms of NRMSE too. Although TimeGrad was constantly the closest competitor baseline in terms of NACRPS, this is not the case here. For the smallest patch setting, DiffSTG was the best baseline in terms of NRMSE, which was outperformed by Diff-Flood by $9.02\%$. For the next two patch settings, DCRNN was the best baseline. In these cases, Diff-Flood outperformed DCRNN by $6.13\%$ and $20.21\%$, respectively. For the settings with $K=20$, DiffSTG and DCRNN both required extensive GPU memory due to poor scalability and they failed to train. In these settings, Diff-Flood outperformed TimeGrad with a margin of around $50-60\%$.
These experiments highlight that existing spatiotemporal forecasting methods (i.e., DiffSTG and DCRNN) have moderate performance but poor computation scalability, making them infeasible to deploy in large spatiotemporal settings. On the other hand, temporal forecasting methods do not have good performance scalability, making them inadequate for large-scale spatiotemporal forecasting tasks. Diff-Flood has both computation and performance scalability, making it suitable for large-scale spatiotemporal forecasting tasks.
Finally, the number of parameters of the baseline models grows almost exponentially with increasing patch sizes ($D$), as observed from Figure 3. To investigate the reason, we analyzed the implementation of TimeGrad as an example and found that it uses a fully connected layer proportional to the number of variables. On the other hand, our model utilizes a UNet architecture that creates a low-dimensional latent representation of the patches. It then performs expensive operations (e.g., feed-forward, cross-attention) on the lower-dimensional latent space, allowing it to scale quite well with the number of variables.
6.4 Ablation Study
Figure 4: Bar-plots showing performance comparison of the four configurations of Diff-Flood on our test dataset of configuration $64\times 64\times 10$, in terms of two metrics. The height of the bars represent mean of the metric and the vertical black lines represent standard deviation. Lower metric value is better. Our proposed architecture of Diff-Flood is able to utilize elevation (ELEV) and temporal covariates (COV) individually, resulting in improved performance in terms of both of our two metrics. Best performance is achieved when both are used.
In this section, we perform an ablation study of the Diff-Flood method to investigate the effects of utilizing elevation matrix and temporal covariates, in addition to the inundation history. Specifically, given inundation history (INUN: ${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$), elevation (ELEV: ${}_{k}\mathbf{s}$), and temporal covariates (COV: $\mathbf{z}_{t_{0}:t-1}$), we define four configurations based on how the context embedding is computed:
  1. 1.
    INUN: Only the historical inundation (${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$) used.
  2. 2.
    INUN, ELEV: We use historical inundation data ${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$ and the elevation matrix ${}_{k}\mathbf{s}$.
  3. 3.
    INUN, COV: We use historical inundation data ${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$ and the temporal covariates $\mathbf{z}_{t_{0}:t-1}$.
  4. 4.
    INUN, ELEV, COV: We include all three components.
We ran five experiment runs for each configuration, where we trained and tested Diff-Flood on the Tidewatch dataset with $K=10$ and $D=64$. Figure 4 shows the performance of the four configurations in-terms of NACRPS and NRMSE. We observe that, configuration (INUN, COV) and (INUN, ELEV) are both better than the configuration (INUN) in terms of the two metrics. This implies that Diff-Flood is able to utilize both elevation and temporal covariates as context effectively. We also observe that, including the temporal covariates results in a greater improvement ($43\%,41\%$ improvement over INUN in terms of NACRPS and NRMSE, respectively) than including spatial context ($8\%,4\%$). However, best performance is achieved when both are utilized ($48\%,47\%$ improvement over INUN).
7 Discussion and Conclusion
Once Diff-Flood is trained, we can use it to answer inundation-scenario related queries relevant to policymakers. Example queries are as follows:
Figure 5: (left) A school area in the Eastern Shore of Virginia. The red cells and four orange patches overlap with the school area. A possible query is: what is the probability that the flooding level in the school area will be above $d$ units within next $T$ hours? (right) A route coming out from the school. The red cells and the eight orange patches overlap with the route. A possible query is: what is the probability that the route coming out of the school will not be flooded in the next $T$ hours?
  • Query 1: What is the probability that the flooding level in an area $A$ will be above $d$ units within next $T$ hours? (Figure 5, left)
  • Query 2: Given a route, what is the probability that the route will not be flooded in the next $T$ hours? (Figure 5, right)
Such queries are of special interest to city officials; answer to these queries will help to better prepare and respond to emergencies. The key idea for answering such queries is that, we have a sampling method in the form a trained Diff-Flood model and we can use it to sample inundation scenarios from the distribution of possible scenarios. We use the sampled scenarios to calculate our desired probability value. In the supplementary materials, we provide a detailed algorithm for answering such queries using a trained Diff-Flood model.
Although Diff-Flood is developed to solve the task of coastal inundation forecasting, it is generalizable enough to be extended to other spatiotemporal forecasting tasks (e.g., occupancy detection, traffic forecasting). Diff-Flood supports conditioning on time-invariant spatial features, as well as, on temporal co-variates. In our future work, we plan to evaluate its performance in diverse spatiotemporal forecasting tasks.
To conclude, we have presented a probabilistic spatiotemporal forecasting method, Diff-Flood that takes both spatial and temporal context into account during its training and forecasting phase. We have performed a case study on the coastal inundation forecasting task for the Eastern Shore of Virginia. Through our experiments, we show that Diff-Flood outperforms existing forecasting methods in predicting coastal inundation. Additionally, the architecture of Diff-Flood allows it to scale to large forecasting tasks at fine spatial resolution, compared to current methods.
Appendix A Extended Related Work
Hydrology Models Traditionally, inundation has been simulated through hydro-dynamical physics-based 1D/2D dual drainage models. By simulating water flow in water bodies and coupling them with 1D or 2D hydrodynamic models, these models can predict inundation in surfaces. For example, the 2D model SOBEK (Deltares, 2014) consists of a 1D network with information about the initial water volume of rivers and 2D rectangular computational cells. By coupling these two parts, the model uses the Saint-Venant flow equation and 2D shallow water equation to calculate discharge between cells and determine the water level at each cell of the Digital Elevation Model (DEM). TELEMAC-2D (Vu et al., 2015) also solves 2D shallow water equation to simulate water flow. However, its primary purpose is to understand flow through water bodies. Therefore, this makes this model more suitable for coastal regions, contrary to SOBEK which is more suitable for urban areas. Polymorphic models like SCHISM (Zhang et al., 2016) can capture water flow at very high resolution, down to even a few meters due to their hybrid grinding system and localized coordinate system. Despite these models’ ability to model inundation accurately at fine resolution, the computation time associated with them makes them infeasible for real-time forecasting.
Regression-based Models Early approaches in finding surrogates for hydrology models involved applying various regression-based methods. (Chang et al., 2010) first applied K-means clustering to pre-process data to find inundation control points. The authors then applied linear regression or neural networks to predict inundation for grids depending on whether they are linearly correlated with the control points or not. (Bermúdez et al., 2019) explored non-parametric regression approach based on least square support vector machine (SVM) as surrogate for 2D hydrodynamic models. (Lee and Kim, 2021) created an offline rainfall-runoff-inundation database by simulating hydrology models over multiple scenarios. Then, they trained a logistic regression-based approach based on the database to estimate the parameters to predict the risk of flooding for each region. Their goal was then to use the trained regressor to generate near real-time prediction of inundation probability given the runoff to the grid from simulated and realistic rainfall scenario.
Deep Learning (DL) Models Many DL models have been explored as possible surrogates for Physics Based Models (PBM). One research work (Pan et al., 2011) applied a combined approach consisting of principal component analysis (PCA) and a feed-forward neural network. Their approach involved inundation prediction at various locations in the next timestep by looking at a fixed window of rainfall history of $N$ different rainfall gauges. First, they applied PCA on these various rainfall gauge histories to reduce the dimensionality before using them as input to the feed-forward neural network. Since such neural networks are expected to fail where data availability is limited, another work (Xie et al., 2021) explored the potential of a hybrid feed-forward neural network-based approach where they used two block-based neural networks along with the traditional point-based neural network and found superiority of the hybrid approach over other in the data-sparse region. Consequently, convolutional neural network (CNN) was explored (Kabir et al., 2020) as a potential surrogate. After their model was trained to predict the output simulated from a 2D hydrology model, they used it to simulate two real-event flood scenarios in a region of UK. Since these models are not able to capture long-term temporal dependencies, researchers have recently explored (Roy et al., 2023) seq2seq LSTM model as surrogates to hydrology models. Such model was also applied in forecasting for multi-step future horizons.
Diffusion Model TimeGrad (Rasul et al., 2021) was the pioneering work that proposed the use of diffusion models for time-series forecasting tasks. They used temporal co-variates and embedded historical data as conditions for the diffusion model to denoise the prediction for the next time step. CSDI Tashiro et al. (2021) does the same but instead of encoding with a Recurrent Neural Network (RNN) as done in TimeGrad, they leverage a transformer-based architecture. SSSDS4 Alcaraz and Strodthoff (2022) uses a S4 model Gu et al. (2021) to capture the encoding which is particularly useful in capturing long-term dependencies. Also, instead of performing diffusion along both temporal and input channel dimensions, they perform diffusion only along the temporal dimension. Different from these, $D^{3}$ (Li et al., 2022) uses coupled diffusion process and bidirectional variational autoencoder to better forecast in longer horizon and improve interpretability. The State-of-the-Art Diffusion-based model DiffSTG (Wen et al., 2023) devised a modified architecture for the reverse process of diffusion model combining UNet and GNN to capture temporal and spatial dependencies, respectively. However, both these methods scale poorly with the growth in number of variables.
Appendix B Denoising Diffusion Model
In this section, we provide an overview of denoising diffusion models.
[…]
$q(\mathbf{x}^{1:N}|\mathbf{x}^{0})=\prod_{n=1}^{N}q(\mathbf{x}^{n}|\mathbf{x}^{n-1})$
is fixed to a Markov chain (often referred to as the forward process), and $(iii)$ the structure of the latent encoder at each hierarchical step is pre-defined as a linear Gaussian model as follows:
$q(\mathbf{x}^{n}|\mathbf{x}^{n-1})=\mathcal{N}(\mathbf{x}^{n};\sqrt{1-\beta_{n}}\mathbf{x}^{n-1},\beta_{n}\mathbf{I}).$
Additionally, the Gaussian parameters ($\beta_{1},\beta_{2},...,\beta_{n})$ are chosen in such a way that we have: $q(\mathbf{x}^{N})=\mathcal{N}(\mathbf{x}^{N};{\mathbf{0}},\mathbf{I})$.
The joint distribution $p_{\theta}(\mathbf{x}^{0:N})$ is called the reverse process. It is defined as a Markov chain with learned Gaussian transitions beginning at $p(\mathbf{x}^{N})$.
$p_{\theta}(\mathbf{x}^{0:N})=p(\mathbf{x}^{N})\prod_{n=1}^{N}p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n}).$
Each subsequent transition is parameterized as follows:
$p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})=\mathcal{N}(\mathbf{x}^{n-1};\mu_{\theta}(\mathbf{x}^{n},n),\Sigma_{\theta}(\mathbf{x}^{n},n)\mathbf{I})$ (20)
Both $\mu_{\theta}:\mathbb{R}^{D}\times\mathbb{N}\rightarrow\mathbb{R}^{D}$ and $\Sigma_{\theta}:\mathbb{R}^{D}\times\mathbb{N}\rightarrow\mathbb{R}^{+}$ take two inputs: $\mathbf{x}^{n}\in\mathbb{R}^{D}$ and the noise step $n\in\mathbb{N}$. The parameters $\theta$ are learned by fitting the model to the data distribution $q_{\chi}(\mathbf{x}^{0})$. This is done by minimizing the negative log-likelihood via a variational bound:
[…]
$=\log\mathbb{E}_{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}\frac{p_{\theta}(\mathbf{x}^{0:N})}{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}$
$\geq\mathbb{E}_{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}\log\frac{p_{\theta}(\mathbf{x}^{0:N})}{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}$ (21)
$\implies-\log p_{\theta}(\mathbf{x}^{0})$ $\leq\mathbb{E}_{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}\log\frac{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}{p_{\theta}(\mathbf{x}^{0:N})}$ (22)
Here, we get (21) by applying Jensen’s inequality. We indirectly minimize the negative log-likelihood of the data, i.e., $-\log p_{\theta}(\mathbf{x}^{0})$, by minimizing the right hand side of (22) . We can show that it is equal to:
$\mathbb{E}_{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}\left[-\log p_{\theta}(\mathbf{x}^{N})+\sum_{n=1}^{N}\log\frac{q(\mathbf{x}^{n}|\mathbf{x}^{n-1})}{p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})}\right]$ (23)
Ho et al. (2020) showed that the forward process has the following property: we can sample $\mathbf{x}^{n}$ using $\mathbf{x}^{0}$ at any noise step $n$ in closed form . Let, $\alpha_{n}=1-\beta_{n}$, and $\bar{\alpha}_{n}=\prod_{i=1}^{n}\alpha_{i}$. Then, we have:
$q(\mathbf{x}^{n}|\mathbf{x}^{0})=\mathcal{N}\left(\mathbf{x}^{n};\sqrt{\bar{\alpha}_{n}}\mathbf{x}^{0},(1-\bar{\alpha}_{n})\mathbf{I}\right).$ (24)
The authors then showed that (23) can be written as KL-divergence between Gaussian distributions:
$-\mathbb{E}_{q(\mathbf{x}^{1}|\mathbf{x}^{0})}\log p_{\theta}(\mathbf{x}^{0}|\mathbf{x}^{1})+D_{KL}\left(q(\mathbf{x}^{N}|\mathbf{x}^{0})\ ||\ p_{\theta}(\mathbf{x}^{N})\right)$
$+\sum_{n=2}^{N}\mathbb{E}_{q(\mathbf{x}^{n}|\mathbf{x}^{0})}\left[D_{KL}(q(\mathbf{x}^{n-1}|\mathbf{x}^{n},\mathbf{x}^{0})\ ||\ p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n}))\right]$ (25)
Here, the authors conditioned the forward process posterior on $\mathbf{x}^{0}$, i.e. $q(\mathbf{x}^{n-1}|\mathbf{x}^{n},\mathbf{x}^{0})$, because then it becomes tractable as follows:
$q(\mathbf{x}^{n-1}|\mathbf{x}^{n},\mathbf{x}^{0})=\mathcal{N}(\mathbf{x}^{n-1};\tilde{\mu}_{n}(\mathbf{x}^{n},\mathbf{x}^{0}),\tilde{\beta}_{n}\mathbf{I})$ (26)
where,
$\tilde{\mu}_{n}(\mathbf{x}^{n},\mathbf{x}^{0})$ $=\frac{\sqrt{\alpha_{n}}(1-\bar{\alpha}_{n-1})}{1-\bar{\alpha_{n}}}\mathbf{x}^{n}+\frac{\sqrt{\bar{\alpha}_{n-1}}(1-\alpha_{n})}{1-\bar{\alpha_{n}}}\mathbf{x}^{0}$ (27)
$\tilde{\beta}_{n}$ $=\frac{1-\bar{\alpha}_{n-1}}{1-\bar{\alpha_{n}}}\beta_{n}$ (28)
Since, the forward process posterior (26) and the backward transition (20) are both Gaussian, we can write the KL-divergence between them as follows:
$D_{KL}\left(q(\mathbf{x}^{n-1}|\mathbf{x}^{n},\mathbf{x}^{0})\ ||\ p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})\right)$
$=\frac{1}{2\tilde{\beta}_{n}}||\tilde{\mu}_{n}(\mathbf{x}^{n},\mathbf{x}^{0})-\mu_{\theta}(\mathbf{x}^{n},n)||_{2}^{2}+C$ (29)
Here, $C$ is a constant independent of $\theta$, and we assume that
$\Sigma_{\theta}=\tilde{\beta_{n}}=\frac{1-\bar{\alpha}_{n-1}}{1-\bar{\alpha_{n}}}\beta_{n}$ (30)
As $\mu_{\theta}(\mathbf{x}^{n},n)$ is conditioned on $\mathbf{x}^{n}$, similar to $\tilde{\mu}_{n}(\mathbf{x}^{n},\mathbf{x}^{0})$ in (27), we can give it the form:
$\mu_{\theta}(\mathbf{x}^{n},n)=\frac{\sqrt{\alpha_{n}}(1-\bar{\alpha}_{n-1})}{1-\bar{\alpha_{n}}}\mathbf{x}^{n}+\frac{\sqrt{\bar{\alpha}_{n-1}}\ \beta_{n}}{1-\bar{\alpha_{n}}}\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)$ (31)
Here, $\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)$ is a neural network that predicts $\mathbf{x}^{0}$ from noisy vector $\mathbf{x}^{n}$ and noise step $n$. With this parameterization, we can show that:
$D_{KL}\left(q(\mathbf{x}^{n-1}|\mathbf{x}^{n},\mathbf{x}^{0})\ ||\ p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})\right)$
$=\frac{1}{2\tilde{\beta}_{n}}\frac{\bar{\alpha}_{n-1}\ \beta_{n}^{2}}{(1-\bar{\alpha}_{n})^{2}}||\mathbf{x}^{0}-\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)||_{2}^{2}$ (32)
Therefore, optimizing a diffusion model can be done by training a neural network to predict the original ground truth vector from an arbitrarily noisy version of it. To approximately minimize the summation term in the objective (25) across all noise steps $n$, we can minimize the expectation over all noise steps as follows:
$\mathop{\mathrm{argmin}}_{\theta}\ \mathbb{E}_{n\sim U[2,N]}\ \mathbb{E}_{q(\mathbf{x}^{n}|\mathbf{x}^{0})}\frac{1}{2\tilde{\beta}_{n}}\frac{\bar{\alpha}_{n-1}\beta_{n}^{2}}{(1-\bar{\alpha}_{n})^{2}}||\mathbf{x}^{0}-\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)||_{2}^{2}$ (33)
This can be done by using stochastic samples over noise steps. Once trained, to sample from the reverse process $\mathbf{x}^{n-1}\sim p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})$, we first sample $\mathbf{x}^{N}$ from $\mathcal{N}({\mathbf{0}},\mathbf{I})$. Then, we compute:
$\mathbf{x}^{n-1}=\mu_{\theta}(\mathbf{x}^{n},n)+\sqrt{\Sigma_{\theta}}\ \mathbf{v}$ (34)
Here, $\mathbf{v}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$ for $n\in[2,N]$ and $\mathbf{v}={\mathbf{0}}$ when $n=1$. $\mu_{\theta}(\mathbf{x}^{n},n)$ and $\Sigma_{\theta}$ are computed using (31) and (30), respectively.
Appendix C Inundation Scenario Queries
Once Diff-Flood is trained, we can use it to answer inundation scenario-based queries relevant to policymakers. In this Section, we discuss a few of such queries.
Figure 6: (left) A school area in the Eastern Shore of Virginia. The red cells and four orange patches overlap with the school area. A possible query is: what is the probability that the flooding level in the school area will be above $d$ units within next $T$ hours? (right) A route coming out from the school. The red cells and the eight orange patches overlap with the route. A possible query is: what is the probability that the route coming out of the school will not be flooded in the next $T$ hours?
Input: Query area $A$ as a Polygon, Flooding level threshold $d$, Prediction Length $T$, Trained Diff-Flood Model.
1 $C_{A}\leftarrow$ Set of grid cells that overlap with $A$.
2 $\mathcal{P}_{A}\leftarrow$ Set of patches where each patch $P\in\mathcal{P}_{A}$ contains at-least one cell $c\in C_{A}$.
3 for each patch $P\in\mathcal{P}_{A}$ do
4       Generate $M$ samples for patch $P$ with prediction length $T$ (using Algorithm 2 in main paper).
5       $num\_scenarios(P_{\leq d})\leftarrow$ Number of samples where each cell $c\in P\cap C_{A}$ has flooding level $\leq d$.
6       $probability(P_{\leq d})\leftarrow num\_scenarios(P_{\leq d})/M$
[…]
return $p$.
Algorithm 3 Calculate the probability that the flooding level in an area $A$ will be $\leq d$ units in the next $T$ hours.
  • Query 1: What is the probability that the flooding level in an area $A$ will be above $d$ units within next $T$ hours?
    We use Algorithm 3 as a sub-routine to answer this query. First, we identify the cells that overlap with $A$ (red cells in Figure 6 left). Let $C_{A}$ denote the set of these cells. If at-least one cell $c\in C_{A}$ has a flooding level above $d$ units in the next $T$ hours, we say area $A$ has a flooding level above $d$ units.
    In Algorithm 3 line 3, we identify the patches that contain at-least one cell $c\in C_{A}$ (four orange patches in Figure 6, left). Let the set of such patches be denoted by $\mathcal{P}_{A}$. For each patch $P\in\mathcal{P}_{A}$, we generate $M$ number of samples (using Algorithm 2 in main paper) with prediction length of $T$ hours (line 3). Then, we calculate the number of samples where all cells $c\in P\cap C_{A}$ have a inundation level $\leq d$ units (line 3). We denote it by $num\_scenarios(P_{\leq d})$. Subsequently, we calculate the probability that all the cells $c\in P\cap C_{A}$ have flooding level $\leq d$ units to be $probability(P_{\leq d})=num\_scenarios(P_{\leq d})/M$ (line 3). We assume the inundation levels in different patches to be independent. Therefore, the probability that all cells $c\in C_{A}$ have flooding level $\leq d$ units is calculated to be $p=\prod_{P\in\mathcal{P}_{A}}probability(P_{\leq d})$ (line 3).
    Since we require the probability of flooding level being above $d$ units in area $A$, our desired probability value is $(1-p)$.
  • Query 2: Given an evacuation route, what is the probability that the route will not be flooded in the next $T$ hours?
    An evacuation route is a sequence of roads that are used for evacuating out of an area. We represent each road using a polygon. Then, a route is also a polygon, which is the union of all of its member road polygons. Let the route polygon be denoted by $A$. Then, we use Algorithm 3 with $d=0$, to calculate our desired probability value.
It is important to note that, we can have more complex queries. For instance, a more refined version of Query 1 can be: what is the probability that at-least half of the region $A$ will have flooding level above $d$ units in the next $T$ hours? The key idea for answering such queries is that, we have a sampling method in the form a trained Diff-Flood model and we can use it to sample inundation scenarios from the distribution of possible scenarios. We use the sampled scenarios to calculate our desired probability value.
Appendix D Hyper-parameters
Hyper-parameter Value
Learning rate 0.001
Learning rate update step size 5 epochs
Learning rate update factor 0.5
Number of epochs 30
Training batch size 32
Validation and test batch size 4
Context length 12 timesteps
Training prediction length 1 timestep
Validation and test prediction length 12 timesteps
Number of diffusion steps 20
Diffusion $\beta_{min}$ $10^{-4}$
Diffusion $\beta_{max}$ 1
NO. of scenarios generated for validation 2
NO. of scenarios generated for test 8
Table 2: Shared hyper-parameters for all of our patch configurations in Table 1 of main paper.
Hyper-parameter $D=16$ $D=64$ $D=80$ $D=96$
Number of Convolution Blocks 1 3 3 3
Convolution block number of Channels [8] [16, 32, 64] [16, 32, 64] [16, 32, 64]
Context embedding dimension 32 32 64 96
UNet Number of Down
(and Up) blocks
4 4 4 4
Number of ResNet layers
per UNet block
2 2 2 2
UNet Down Blocks:
Number of Channels
[8, 16, 32, 32] [16, 32, 32, 64] [16, 32, 32, 64] [16, 32, 32, 64]
UNet Number of Cross-
attention Down / Up Blocks
2 2 2 2
Group Normalization:
Number of Groups
8 8 8 8
Table 3: Hyper-parameters for different patch configurations
Appendix E Patch Visualization
Figure 7: Visualization of inundation on $10$, $64\times 64$ patches in the first $24$ hours of training data. The darker the shade of blue, the higher the inundation value.

### v2

@@
-Denoising Diffusion Probabilistic Models for Coastal Inundation Forecasting
+Towards High Resolution Probabilistic Coastal Inundation Forecasting from Sparse Observations
 Abstract
-Coastal flooding poses significant risks to communities, necessitating fast and accurate forecasting methods to mitigate potential damage.
-To approach this problem, we present Diff-Flood, a probabilistic spatiotemporal forecasting method designed based on denoising diffusion models.
-Diff-Flood predicts inundation level at a location by taking both spatial and temporal context into account.
-It utilizes inundation levels at neighboring locations and digital elevation data as spatial context.
-Inundation history from a context time window, together with additional co-variates are used as temporal context.
-Convolutional neural networks and cross-attention mechanism are then employed to capture the spatiotemporal dynamics in the data.
-We trained and tested Diff-Flood on coastal inundation data from the Eastern Shore of Virginia, a region highly impacted by coastal flooding.
-Our results show that, Diff-Flood outperforms existing forecasting methods in terms of prediction performance ($6\%\text{ to }64\%$ improvement in terms of two performance metrics) and scalability.
-1 Introduction
-The projection in sea level rise have highlighted the increasing vulnerability of many coastal regions to inundation (Wuebbles et al., 2017).
-Among these regions, the East Coast in US is a region of particular policy concern due to its geography and low-lying topography (Ezer and Atkinson, 2014).
-This region, in recent periods, has been experiencing frequent inundation events  (Ezer, 2020), causing disruptions in daily life (e.g.
-traffic, economic) (Jacobs et al., 2018).
-In the future, this situation will only keep declining due to the combined effect of climate change, sea level rise, and urbanization (Swain et al., 2020).
-Therefore, forecasting coastal inundation is of utmost importance from a policy standpoint for undertaking risk prevention measurement or safe evacuation of the residents in danger  (Yang et al., 2019).
-Figure 1: Flooding on Quinby Bridge Road, Accomack County, VA on April 3, 2024 (Data source: Tidewatch Maps (Tidewatch, 2025).)
-The problem of coastal inundation forecasting came to our attention during conversations with staff at the City of Virginia Beach.
+Coastal flooding poses increasing threats to communities worldwide, necessitating accurate and hyper-local inundation forecasting for effective emergency response.
+However, real-world deployment of forecasting systems is often constrained by sparse sensor networks, where only a limited subset of locations may have sensors due to budget constraints.
+To approach this challenge, we present Diff-Sparse, a masked conditional diffusion model designed for probabilistic coastal inundation forecasting from sparse sensor observations.
+Diff-Sparse primarily utilizes the inundation history of a location and its neighboring locations from a context time window as spatiotemporal context.
+The fundamental challenge of spatiotemporal prediction based on sparse observations in the context window is addressed by introducing a novel masking strategy during training.
+Digital elevation data and temporal co-variates are utilized as additional spatial and temporal contexts, respectively.
+A convolutional neural network and a conditional UNet architecture with cross-attention mechanism are employed to capture the spatiotemporal dynamics in the data.
+We trained and tested Diff-Sparse on coastal inundation data from the Eastern Shore of Virginia and systematically assessed the performance of Diff-Sparse across different sparsity levels ($0\%,50\%,95\%$ missing observations).
+Our experiment results show that Diff-Sparse achieves upto $62\%$ improvement in terms of two forecasting performance metrics compared to existing methods, at $95\%$ sparsity level.
+Moreover, our ablation studies reveal that digital elevation data becomes more useful at high sparsity levels compared to temporal co-variates.
+Codehttps://github.com/KAI10/Diff-Sparse
+Introduction
+The projection in sea level rise have highlighted the increasing vulnerability of many coastal regions to inundation (wuebbles2017climate).
+Among these regions, the East Coast in US is a region of particular policy concern due to its geography and low-lying topography (ezer2014accelerated).
+In recent periods, this region has been experiencing frequent inundation events  (ezer2020analysis), causing disruptions in daily life (e.g., traffic, economic) (jacobs2018recent).
+In future, this situation will only continue to decline due to the combined effect of climate change, sea level rise, and urbanization (swain2020increased).
+Therefore, forecasting coastal inundation is of utmost importance from a policy standpoint for undertaking risk prevention measurement and safe evacuation of residents during emergencies  (yang2019incorporating).
+The problem of forecasting coastal inundation from sparse observation was brought to our attention through conversations with staff at the City of Virginia Beach.
 The region faces challenges with coastal flooding which causes slow-down and/or road closures; disrupting the day-to-day activities of the inhabitants.
-A fast and high-resolution flood forecasting model would enable city officials to identify localized inundation risks, resulting in more effective preparedness and response.
-In this paper, we approach the problem of designing such a model.
-It is a joint-work of a multidisciplinary team, including computer and data scientists, academic experts with expertise in hydro-dynamic models, members of established coastal resources management centers, and officials at the City of Virginia Beach.
+A fast and high-resolution flood forecasting model would enable city officials to identify localized flood risks, resulting in more effective preparedness and response.
+In order to effectively identify inundated locations, policymakers deploy sensors at various locations (tao2024review).
+These sensors are responsible for providing periodical inundation reports and other variables of interest in the surrounding area.
+However, there are several challenges associated with these sensor placements.
+First, due to budget and resource constraints, the number of sensors that can be deployed is limited (saad2024scalable; karagulian2019review).
+Second, the readings from the sensor are often associated with noise due to mechanical, environmental or calibration issues (barrenetxea2008assessing; mydlarz2024floodnet).
+Third, due to the difficulty associated with finding an optimal location assignment of spatial sensors (krause2008near) or changes carried out by the stakeholders, the locations of these sensors may change from time to time (garg2012learning).
+As a result, a forecasting method that relies on sensor data but does not account for the sparsity and uncertain nature of sensor placements will not be reliable (tao2024review).
 The traditional method for predicting inundation involves hydro-dynamical, physics-based models (PBM).
-Commercially available (Syme, 2001; Deltares, 2014) (e.g., SOBEK, TUFLOW) or open-sourced (Zhang et al., 2016) (e.g, SCHISM) hydro-dynamical models can simulate inundation accurately at a high resolution.
-However, the computation time associated with these PBMs makes them challenging for real-time forecasting (Guo and others, 2021).
-Although researchers have investigated Deep-Learning (DL) based time-series prediction methods as surrogates to these PBMs (Roy et al., 2023), these models are not suitable for accommodating spatial contexts.
-Existing spatiotemporal forecasting methods, on the other hand, struggle to scale in high-resolution forecasting tasks.
-Motivated by the above observations, we explore the Diffusion Model (Sohl-Dickstein et al., 2015) in the context of high-resolution coastal inundation forecasting task.
-The primary reason for investigating diffusion models is its ability to utilize conditioning contexts of various dimensions (Rombach et al., 2022; Zhang and others, 2023) and its potential to model physical properties (Ahmad and others, 2024; Amram and Pedro, 2023).
-Originally developed for image generation, diffusion models have been adapted for time-series forecasting, achieving state-of-the-art performance, as demonstrated by Rasul et al.
-(2021).
-However, their approach does not incorporate spatial context.
-More recently,  (Wen et al., 2023) proposed a probabilistic spatiotemporal graph forecasting method based on denoising diffusion models.
-While promising, such methods face scalability challenges when applied to large-scale/high-resolution inundation forecasting tasks.
-Using the coastal inundation of the Eastern Shore as case study, we address these challenges by making the following contributions in this paper.
-  • We formulate the problem of coastal inundation forecasting under the spatiotemporal forecasting framework and introduce Diff-Flood as a novel method for approaching the problem.
-It can utilize both spatial and temporal context to generate spatiotemporal forecasts at fine spatial resolution.
-  • Extensive evaluation of our method against several baseline methods on inundation data of the Virginia Eastern Shore underscores the superiority of our proposed method and its aspect of being used as a decision-making tool for policymakers.
-Specifically, we observed $6\%$ to $64\%$ improvement in prediction performance, in terms of two performance metrics (Normalized Average Continuous Ranked Probability Score, and Normalized Root Mean Square Error).
-An ablation study of our model is performed to highlight the importance of its components.
-  • Furthermore, we demonstrate the scalability of our model compared to other baselines in large settings, underscoring its utility in real-time forecasting in large-scale scenarios.
-It is worth highlighting that our model can be used to answer inundation scenario-based queries relevant to policymakers.
-For example, queries of the form: “What is the probability that a particular region will get inundated by 1 feet within the next 12 hours?”, can be answered by using our model through probabilistic forecasting, see section 7.
-2 Related Work
+Commercially available (syme2001tuflow; deltares2014sobek) (e.g., SOBEK, TUFLOW) or open-sourced (zhang2016seamless) (e.g, SCHISM) hydro-dynamical models can simulate inundation accurately at a high resolution.
+However, the computation time associated with these PBMs makes them challenging for real-time forecasting (guo2021urban).
+Researchers have investigated the use of Deep-Learning (DL) based prediction methods as surrogates to these PBMs (roy2023application).
+However, these models often struggle to: (i) scale in large-scale high-resolution forecasting tasks under sparse availability of data, and (ii) effectively use additional spatiotemporal contexts.
+Therefore, a surrogate model to PBMs that can make accurate real-time predictions based on sparse sensor observations, will have high relevance to policymakers.
+Motivated by the above observations, we explore Diffusion Models (sohl2015deep) for high-resolution coastal inundation forecasting task.
+The reason is two-fold: their potential to model physical properties (ahmad2024calobench; amram2023denoising), and their ability to utilize conditioning contexts of various dimensions (Rombach_2022_CVPR; zhang2023adding) that can improve prediction quality in the presence of sparse observations.
+Although diffusion models were originally developed for image generation, they have been adapted for time-series and spatiotemporal forecasting (rasul2021autoregressive; wen2023diffstg).
+However, the existing methods struggle in terms of computational scalability and performance, under sparse availability of data.
+Using the coastal inundation of the Virginia Eastern Shore as case study, we address these challenges by making the following contributions:
+  • We formulate the problem of probabilistic inundation forecasting from sparse observations as a conditional distribution function learning problem and present Diff-Sparse, a scalable model to learn this function.
+Diff-Sparse is trained using data from physics-based hydro-dynamic model; however, during inference it makes probabilistic forecasts based on sparse sensor observations and additional spatiotemporal contexts (e.g., elevation, temporal co-variates).
+  • We introduce a novel masking strategy to train Diff-Sparse which enables it to make forecasts from sparse observations.
+Our robust training strategy enables Diff-Sparse to make accurate spatiotemporal forecasts from different spatial placement configurations of the sensors, without requiring training the model from scratch.
+  • Extensive evaluation of our method against several baseline methods under sparse observation underscores the superiority of our method at large-scale high resolution inundation forecasting tasks.
+Specifically, we observe upto $62\%$ improvement in prediction performance over existing methods in terms of two predictive performance metrics.
+  • We perform extensive ablation studies to highlight the utility of different context data (e.g., sensor data, elevation, temporal co-variates) at varying sparsity levels.
+Related Work
 We present an abridged version of the most relevant literature in this section. The supplementary material of this paper covers the literature more extensively.
-Hydrology Models: Traditionally, various Physics Based hydrodynamic Models (PBM) have been employed to simulate inundation.
-SOBEK (Deltares, 2014) is a 2D model that solves a Saint-Venant flow equation and shallow water equation to determine the water level of rectangular grids.
-TELEMAC-2D (Vu et al., 2015) uses a similar method but is more suitable for coastal areas.
-Polymorphic models like SCHISM (Zhang et al., 2016) can capture water flow at very high resolution.
-However, the computation time associated with solving these complex equations at high spatial resolution makes these PBMs infeasible for real-time forecasting.
-Deep Learning: DL methods are a popular choice in surrogating hydrodynamic models.
-Regression-based methods (Zahura and Goodall, 2022), recurrent neural networks (Roy et al., 2023) and hybrid ML-based methods (Zanchetta and Coulibaly, 2022) are some of the methods that have been attempted as possible surrogates.
-However, these models cannot inherently take spatial context into account.
-The DCRNN method proposed by (Li et al., 2017) for traffic forecasting is the state-of-the-art DL method for spatiotemporal point forecasting, where bidirectional random walk is used to capture spatial dependency and Seq2Seq architecture is used to capture temporal dependency.
+Physcis-based Models: Traditionally, various Physics-Based hydrodynamic Models (PBMs) have been employed to simulate inundation.
+SOBEK (deltares2014sobek) is a two-dimensional model that solves the Saint-Venant flow equations and shallow water equations to determine water levels across rectangular grids.
+Similarly, TELEMAC-2D (vu2015two) employs comparable method specifically tailored for coastal applications.
+Polymorphic models such as SCHISM (zhang2016seamless) captures water flow at very high resolutions.
+However, the substantial computational time associated with solving these complex equations at high spatial resolution renders these PBMs impractical for real-time forecasting applications.
+Deep Learning Researchers have explored Deep Learning (DL) methods for developing surrogates to PBMs.
+Regression-based methods (zahura2022predicting), recurrent neural networks (roy2023application), and hybrid machine learning-based methods (zanchetta2022hybrid) have been emerged as popular choices, among others.
+However, these models inherently struggle to incorporate spatiotemporal contexts effectively.
+Originally developed for traffic forecasting, the Diffusion Convolutional Recurrent Neural Network (DCRNN) method proposed by Li et al.
+li2017diffusion captures some aspect of spatiotemporal contexts by utilizing bidirectional random walks and Sequence-to-Sequence architecture.
+More recently, BayesNF (saad2024scalable) has emerged as the state-of-the-art DL spatiotemporal forecasting model, constructed using a Bayesian Neural Network framework.
+By assigning prior distributions to model parameters and adjusting the posterior based on observations, BayesNF effectively learns to map multivariate spatiotemporal points to continuous real-valued fields.
 Diffusion Model The field of generative modeling has been dominated by the diffusion model for quite some time.
-Although it had primarily been studied for image synthesis (Ho et al., 2020; Dhariwal and Nichol, 2021; Austin et al., 2021), its fascinating ability to use flexible conditioning to guide the generation process has been utilized in other domains like video generation (Yang et al., 2023b), prompt-based image generation (Ramesh et al., 2022), text-to-speech generation (Yang et al., 2023a).
-TimeGrad (Rasul et al., 2021) was the pioneering method that approached the time-series forecasting problem by utilizing a diffusion model.
-Other models like CSDI (Tashiro et al., 2021), DPSD-CSPD (Biloš et al., 2023), TDSTF (Chang et al., 2024) were later introduced.
+Although it had primarily been studied for image synthesis (ho2020denoising; dhariwal2021diffusion; austin2021structured), its fascinating ability to use flexible conditioning to guide the generation process has been utilized in other domains like video generation (yang2023diffusion), prompt-based image generation (ramesh2022hierarchical), text-to-speech generation (yang2023diffsound).
+TimeGrad (rasul2021autoregressive) was the pioneering method that approached the time-series forecasting problem by utilizing a diffusion model.
+Other models like CSDI (tashiro2021csdi), DPSD-CSPD (bilovs2023modeling), TDSTF (Chang2024) were later introduced.
 These models try to predict the noise added during the forward process of the diffusion model ($\epsilon$-parameterization).
 An alternative parameterization is to predict the ground-truth data ($x$-parameterization).
-Different from models mentioned above, $D^{3}$ (Li et al., 2022) uses coupled diffusion process and bidirectional variational autoencoder to better forecast in longer horizon and improve interpretability.
-The State-of-the-Art Diffusion-based model DiffSTG (Wen et al., 2023) devised a modified architecture for the reverse process of diffusion model combining UNet and GNN to capture temporal and spatial dependencies, respectively.
-However, both these methods scale poorly with the growth in number of variables.
-3 Problem Formulation
-Formally, we define the inundation forecasting problem as follows.
-Problem 1.
-[…]
-Problem 2.
-Multi-patch Probabilistic Inundation Forecasting Let $\mathcal{P}=\{P_{1},P_{2},...,P_{K}\}$ denote a set of two-dimensional patches of size ${D\times D}$.
-Let ${}_{k}\mathbf{x}_{t}^{0}\in\mathbb{R}^{D\times D}$ denote the inundation matrix of patch $P_{k}$ at timestep $t\in\{1,\ldots,T\}$.
-Let ${}_{k}\mathbf{s}\in\mathbb{R}^{D\times D}$ and ${}_{k}\mathbf{z}_{t}\in\mathbb{R}^{l},t\in\{1,\ldots,T\}$ denote the elevation matrix and time series of co-variates associated with patch $P_{k}$, respectively.
+Different from models mentioned above, $D^{3}$VAE (li2022generative) uses coupled diffusion process and bidirectional variational autoencoder to better forecast in longer horizon and improve interpretability.
+The State-of-the-Art Diffusion-based model DiffSTG (wen2023diffstg) devised a modified architecture for the reverse process of diffusion model combining UNet and GNN to capture temporal and spatial dependencies, respectively.
+However, both methods scale poorly with the growth of variables.
+Problem Formulation
+Formally, we define the inundation forecasting from sparse observations problem as follows.
+Problem 1
+Multi-patch Probabilistic Inundation Forecasting.
+Let $\mathcal{P}=\{P_{1},P_{2},...,P_{K}\}$ denote a set of two-dimensional rectangular grids each of size $D\times D$.
+We refer to $P_{k}$ as the $k^{th}$ ‘patch’.
+Let $\mathcal{M}=\{\mathbf{M}_{1},\mathbf{M}_{2},...,\mathbf{M}_{K}\}$ denote a set of binary matrices, henceforth referred to as ‘sensor masks’.
+If $\mathbf{M}_{k}(i,j)=1$ then the cell $(i,j)$ of patch $P_{k}$ has a water-level sensor; $\mathbf{M}_{k}(i,j)=0$ means cell $(i,j)$ of patch $P_{k}$ does not have a water-level sensor.
+Let $_{k}\mathbf{y}_{t}^{0}\in\mathbb{R}^{D\times D}$ denote the inundation matrix of patch $P_{k}$ at timestep $t$.
+If $\mathbf{M}_{k}(i,j)=1$ then $_{k}\mathbf{y}_{ij,t}^{0}\in\mathbb{R}$ contains the inundation level on cell $(i,j)$ of patch $P_{k}$ at time step $t$; $i,j\in\left\{1,...,D\right\},t\in\{1,...,T\}$.
+However, for cells $(i,j)$ where $\mathbf{M}_{k}(i,j)=0$, $_{k}\mathbf{y}_{ij,t}^{0},t\in\{1,...,T\}$ has missing value.
+Let $_{k}\mathbf{x}_{t}^{0}\in\mathbb{R}^{D\times D}$ denote the complete inundation matrix of patch $P_{k}$ at timestep $t$, where there are no missing values.
+Let $\mathbf{s}_{k}\in\mathbb{R}^{D\times D}$ and $_{k}\mathbf{z}_{t}\in\mathbb{R}^{l},t\in\{1,\ldots,T\}$ denote the elevation matrix and time series of co-variates associated with patch $P_{k}$, respectively.
 Let $t_{0}:t_{c+l}$ denote the increasing sequence of timesteps $t_{0},t_{1},...,t_{c+l}$}, where $c,l>0$.
-We aim to learn the conditional distribution function $q({}_{k}\mathbf{x}_{t_{c}:t_{c+l}}^{0}|{}_{k}\mathbf{x}_{t_{0}:t_{c-1}}^{0},{}_{k}\mathbf{z}_{t},{}_{k}\mathbf{s})$.
-Note that, both problems are spatio-temporal forecasting problems, since we aim to forecast the inundation level on all cells in a patch together.
-They are also probabilistic forecasting problems, as we want to learn the distribution of inundation values instead of making a single prediction.
-Finally, Problem 2 represents a generalized version of Problem 1.
-Here, we aim to learn a single model that can be utilized to forecast inundation levels on multiple patches/areas.
-The following section presents the basics of a denoising diffusion model used by our method Diff-Flood for learning our target distributions.
-Subsequently, we present details on how we have designed Diff-Flood to capture spatial and temporal dynamics of our inundation values.
-4 Denoising Diffusion Model
-In this section, we provide a brief overview of denoising diffusion models. A more comprehensive view is provided in the supplementary material.
-[…]
+We aim to learn the conditional distribution function $q({}_{k}\mathbf{x}_{t_{c}:t_{c+l}}^{0}|{}_{k}\mathbf{y}_{t_{0}:t_{c-1}}^{0},{}_{k}\mathbf{z}_{t},\mathbf{s}_{k},\mathbf{M}_{k})$.
+Here, we aim to forecast the inundation on all cells in a patch (i.e., $\mathbf{x}$ ) even though in the context inundation history (i.e., $\mathbf{y}$) we only know the inundation levels on cells where sensors are placed.
+Problem 1 is a probabilistic forecasting problem, as we want to learn the distribution of inundation values instead of making a single prediction.
+The following section presents the basics of a denoising diffusion model used by our method Diff-Sparse for learning our target distributions.
+Subsequently, we present details on how we have designed Diff-Sparse to capture spatial and temporal dynamics of inundation values.
+Denoising Diffusion Model
+In this section, we provide a brief overview of denoising diffusion models. A more comprehensive view is provided in the supplementary materials.
+Let $\mathbf{x}^{0}\sim q_{\chi}(\mathbf{x}^{0})$ denote the multivariate training vector from some input space $\chi=\mathbb{R}^{D}$, where $q_{\chi}(\mathbf{x}^{0})$ denotes the true distribution.
+In diffusion probabilistic models, we aim to approximate $q_{\chi}(\mathbf{x}^{0})$ by a probability density function $p_{\theta}(\mathbf{x}^{0})$ where $\theta$ denotes the set of trainable parameters.
+Diffusion probabilistic models (sohl2015deep; luo2022understanding) are a special class of Hierarchical Variational Autoencoders (kingma2016improved; sonderby2016ladder) with the form $p_{\theta}(\mathbf{x}^{0})=\int p_{\theta}(\mathbf{x}^{0:N})d\mathbf{x}^{1:N}$; here $\mathbf{x}^{1},\mathbf{x}^{2},...,\mathbf{x}^{N}$ are latent variables.
+Three key properties of diffusion models are: $(i)$ all of the latent variables are assumed to have the dimension $D$ as $\mathbf{x}^{0}$, $(ii)$ the approximate posterior ,
 $q(\mathbf{x}^{1:N}|\mathbf{x}^{0})=\prod_{n=1}^{N}q(\mathbf{x}^{n}|\mathbf{x}^{n-1})$
-is fixed to a Markov chain (often referred to as the forward process), and $(iii)$ the structure of the latent encoder at each hierarchical step is pre-defined as a linear Gaussian model as follows:
+is fixed to a Markov chain (often referred to as the forward process), and $(iii)$ the structure of the latent encoder at each hierarchical step is pre-defined as a linear Gaussian model:
 $q(\mathbf{x}^{n}|\mathbf{x}^{n-1})=\mathcal{N}(\mathbf{x}^{n};\sqrt{1-\beta_{n}}\mathbf{x}^{n-1},\beta_{n}\mathbf{I}).$
@@
 This is done by minimizing the negative log-likelihood via a variational bound:
-[…]
-$=\log\mathbb{E}_{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}\frac{p_{\theta}(\mathbf{x}^{0:N})}{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}$
-$\geq\mathbb{E}_{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}\log\frac{p_{\theta}(\mathbf{x}^{0:N})}{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}$ (2)
-$\implies-\log p_{\theta}(\mathbf{x}^{0})$ $\leq\mathbb{E}_{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}\log\frac{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}{p_{\theta}(\mathbf{x}^{0:N})}$ (3)
-Here, we get (2) by applying Jensen’s inequality. We indirectly minimize the negative log-likelihood of the data, i.e., $-\log p_{\theta}(\mathbf{x}^{0})$, by minimizing the right hand side of (3).
-Ho et al.
-(2020) showed that the forward process has the following property: we can sample $\mathbf{x}^{n}$ using $\mathbf{x}^{0}$ at any noise step $n$ in closed form .
+$-\log p_{\theta}(\mathbf{x}^{0})\leq\mathbb{E}_{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}\log\frac{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}{p_{\theta}(\mathbf{x}^{0:N})}$ (2)
+We indirectly minimize the negative log-likelihood of the data by minimizing the right hand side of (2).
+(ho2020denoising) showed that the forward process has the following property: we can sample $\mathbf{x}^{n}$ using $\mathbf{x}^{0}$ at any noise step $n$ in closed form .
 Let, $\alpha_{n}=1-\beta_{n}$, and $\bar{\alpha}_{n}=\prod_{i=1}^{n}\alpha_{i}$.
@@
 $q(\mathbf{x}^{n}|\mathbf{x}^{0})=\mathcal{N}\left(\mathbf{x}^{n};\sqrt{\bar{\alpha}_{n}}\mathbf{x}^{0},(1-\bar{\alpha}_{n})\mathbf{I}\right).$ (3)
-Based on this property, Ho et al. (2020) showed that one possible parameterization of (1) is as follows:
+Based on this property, (ho2020denoising) showed that one possible parameterization of (1) is as follows:
 $\mu_{\theta}(\mathbf{x}^{n},n)$ $=\frac{\sqrt{\alpha_{n}}(1-\bar{\alpha}_{n-1})}{1-\bar{\alpha_{n}}}\mathbf{x}^{n}+\frac{\sqrt{\bar{\alpha}_{n-1}}\ \beta_{n}}{1-\bar{\alpha_{n}}}\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)$ (4)
@@
 Here, $\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)$ is a neural network that predicts $\mathbf{x}^{0}$ from noisy vector $\mathbf{x}^{n}$ and noise step $n$.
-Ho et al.
-(2020) then showed that, the right hand side of (3) can be minimized by minimizing:
-$\mathbb{E}_{n,q(\mathbf{x}^{n}|\mathbf{x}^{0})}\ \frac{1}{2\tilde{\beta}_{n}}\frac{\bar{\alpha}_{n-1}\ \beta_{n}^{2}}{(1-\bar{\alpha}_{n})^{2}}||\mathbf{x}^{0}-\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)||_{2}^{2}$ (7)
+(ho2020denoising) then showed that, the right hand side of (2) can be minimized by minimizing:
+[…]
 Therefore, optimizing a diffusion model can be done by training a neural network to predict the original ground truth vector from an arbitrarily noisy version of it.
@@
 Then, we compute:
-$\mathbf{x}^{n-1}=\mu_{\theta}(\mathbf{x}^{n},n)+\sqrt{\Sigma_{\theta}}\ \mathbf{v}$ (8)
-Here, $\mathbf{v}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$ for $n\in[2,N]$, and $\mathbf{v}={\mathbf{0}}$ when $n=1$.
-$\mu_{\theta}(\mathbf{x}^{n},n)$ and $\Sigma_{\theta}$ are computed using (5) and (6), respectively.
-5 Diff-Flood Method
-Diff-Flood is designed for multi-patch probabilistic inundation forecasting, where the inundation values are correlated both spatially and temporally.
-Following our notations in Problem 1 and 2, let ${}_{k}\mathbf{x}_{t}^{0}\in\mathbb{R}^{D\times D}$ denote the inundation matrix of the patch $P_{k}$ at time step $t$.
-Towards our goal of learning the conditional distribution $q_{\chi}({}_{k}\mathbf{x}_{t_{c}:t_{c+l}}^{0}|{}_{k}\mathbf{x}_{t_{0}:t_{c-1}}^{0},{}_{k}\mathbf{z}_{t_{0}:t_{c-1}},{}_{k}\mathbf{s})$, where ${}_{k}\mathbf{s}\in\mathbb{R}^{D\times D}$ is the elevation matrix of patch $P_{k}$, and ${}_{k}\mathbf{z}_{t}$ denotes a time-series of co-variates associated with $P_{k}$, we assume:
-$q_{\chi}({}_{k}\mathbf{x}_{t_{c}:t_{c+l}}^{0}|{}_{k}\mathbf{x}_{t_{0}:t_{c-1}}^{0},{}_{k}\mathbf{z}_{t_{0}:t_{c-1}},{}_{k}\mathbf{s})=$
-$\quad\quad\quad\textstyle\prod_{t=t_{c}}^{t_{c+l}}q_{\chi}({}_{k}\mathbf{x}_{t}^{0}|{}_{k}\mathbf{x}_{t_{0}:t-1}^{0},{}_{k}\mathbf{z}_{t_{0}:t-1},{}_{k}\mathbf{s})$ (9)
-In Diff-Flood, we use a denoising diffusion model (as described in Section B) to learn the conditional distributions on the right hand side of (9).
-We refer to ${}_{k}\mathbf{x}_{t_{0}:t-1}^{0}$, ${}_{k}\mathbf{z}_{t_{0}:t-1}$, and ${}_{k}\mathbf{s}$ as context data.
-Figure 2: Architecture of proposed Diff-Flood Model
-To capture spatial correlation, we first augment the inundation data in the context (i.e.
-${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$) by concatenating the elevation matrix ${}_{k}\mathbf{s}$ to it as an additional channel (Equation 10).
+$\mathbf{x}^{n-1}=\mu_{\theta}(\mathbf{x}^{n},n)+\sqrt{\Sigma_{\theta}}\ \mathbf{u}$ (7)
+Here, $\mathbf{u}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$ for $n\in[2,N]$, and $\mathbf{u}={\mathbf{0}}$ when $n=1$.
+$\mu_{\theta}(\mathbf{x}^{n},n)$, $\Sigma_{\theta}$ are computed using (4) and (5), respectively.
+Diff-Sparse Method
+Diff-Sparse is designed for probabilistic inundation forecasting from sparse observations.
+Following our notations in Problem 1, let $_{k}\mathbf{y}_{t}^{0}\in\mathbb{R}^{D\times D}$ denote the ‘sparse’ inundation matrix of the patch $P_{k}$ at time step $t$, with sensor mask $\mathbf{M}_{k}$.
+Let $_{k}\mathbf{x}_{t}^{0}\in\mathbb{R}^{D\times D}$ denote the complete inundation matrix of patch $P_{k}$ at timestep $t$.
+Towards our goal of learning the conditional distribution $q_{\chi}({}_{k}\mathbf{x}_{t_{c}:t_{c+l}}^{0}|{}_{k}\mathbf{y}_{t_{0}:t_{c-1}}^{0},{}_{k}\mathbf{z}_{t_{0}:t_{c-1}},\mathbf{s}_{k},\mathbf{M}_{k})$, where $\mathbf{s}_{k}\in\mathbb{R}^{D\times D}$ is the elevation matrix of patch $P_{k}$, and $_{k}\mathbf{z}_{t}$ denotes a time-series of co-variates associated with $P_{k}$, we assume:
+$q_{\chi}({}_{k}\mathbf{x}_{t_{c}:t_{c+l}}^{0}|{}_{k}\mathbf{y}_{t_{0}:t_{c-1}}^{0},{}_{k}\mathbf{z}_{t_{0}:t_{c-1}},\mathbf{s}_{k},\mathbf{M}_{k})=$
+$\quad\quad\quad\textstyle\prod_{t=t_{c}}^{t_{c+l}}q_{\chi}({}_{k}\mathbf{x}_{t}^{0}|{}_{k}\mathbf{y}_{t_{0}:t-1}^{0},{}_{k}\mathbf{z}_{t_{0}:t-1},\mathbf{s}_{k},\mathbf{M}_{k})$ (8)
+In Diff-Sparse, we use a denoising diffusion model to learn the conditional distributions on the right hand side of (8).
+We refer to $_{k}\mathbf{y}_{t_{0}:t-1}^{0}$, $_{k}\mathbf{z}_{t_{0}:t-1}$, $\mathbf{s}_{k}$, and $M_{k}$ together as context data.
+Figure 1: Diff-Sparse architecture.
+It has two main components.
+The first component involves context learning, where the sparse context inundation data ($\mathbf{y}_{t_{0}},\mathbf{y}_{t_{1}},...,\mathbf{y}_{t_{c-1}}$), sensor mask ($\mathbf{M}$), elevation data ($\mathbf{s}$) and temporal covariates ($\mathbf{z}_{t_{0}:t_{c-1}}$) are used to generate context embedding $\mathbf{h}_{t_{0}:t_{c-1}}$.
+The second component involves a diffusion model where the corresponding forecast is sampled by conditioning on the context embedding.
+$q(\mathbf{x}_{t_{c}}^{n}|\mathbf{x}_{t_{c}}^{n-1})$ denotes the forward diffusion process; $p_{\theta}(\mathbf{x}_{t_{c}}^{n-1}|\mathbf{x}_{t_{c}}^{n},n,\mathbf{h}_{t_{0}:t_{c-1}})$ denotes the reverse diffusion process, parameterized by $\theta$.
+To capture spatial correlation, we augment the sparse inundation data in the context (i.e., $_{k}\mathbf{y}^{0}_{t_{0}:t-1}$) by concatenating the elevation matrix $\mathbf{s}_{k}$ and the sensor mask $\mathbf{M}_{k}$ to it as additional channels (9).
 The goal is to capture the correlation of elevation and inundation values.
-We then use convolution blocks extract spatial features ${}_{k}\mathbf{f}_{t_{0}:t-1}$ (Equation 11).
-${}_{k}\mathbf{y}^{0}_{t_{0}:t-1}={}_{k}\mathbf{x}^{0}_{t_{0}:t-1}\oplus{}_{k}\mathbf{s}$ (10)
-${}_{k}\mathbf{f}_{t_{0}:t-1}=\textsc{conv}{}_{\theta_{1}}({}_{k}\mathbf{y}^{0}_{t_{0}:t-1})$ (11)
-In addition, since coastal inundation is expected to be correlated with tide cycles, we use the hour of day, and day of month as temporal features $\mathbf{z}_{t}$.
-We use sinusoidal encoding to capture the cyclic nature of these temporal features.
-${}_{k}\mathbf{g}^{0}_{t_{0}:t-1}=\textsc{sinusoid}{}({}_{k}\mathbf{z}_{t_{0}:t-1})$ (12)
+We then use convolution blocks to extract spatial features $_{k}\mathbf{f}_{t_{0}:t-1}$ (10).
+${}_{k}\mathbf{v}_{t_{0}:t-1}={}_{k}\mathbf{y}^{0}_{t_{0}:t-1}\oplus\mathbf{s}_{k}\oplus M_{k}$ (9)
+[…]
+Since coastal inundation is expected to be correlated with tide cycles, we use hour of day, and day of month as temporal features $\mathbf{z}_{t}$.
+We use sinusoidal encoding (vaswani2017attention) to capture the cyclic nature of these temporal features.
+${}_{k}\mathbf{g}_{t_{0}:t-1}=\textsc{sinusoid}{}({}_{k}\mathbf{z}_{t_{0}:t-1})$ (11)
 We concatenate the spatial and temporal features, and then pass it through a linear fully connected layer to get an embedding of the context.
-${}_{k}\mathbf{h}_{t_{0}:t-1}=\textsc{linear}{}_{\theta_{2}}({}_{k}\mathbf{f}_{t_{0}:t-1}\oplus{}_{k}\mathbf{g}_{t_{0}:t-1})$ (13)
-An overview of the architecture of Diff-Flood is provided in Figure 2.
+[…]
+An overview of Diff-Sparse architecture is provided in Figure 1.
 We approximate (8) by the model in (13).
@@
 $\textstyle\prod_{t=t_{c}}^{t_{c+l}}p_{\theta}({}_{k}\mathbf{x}_{t}^{0}|{}_{k}\mathbf{h}_{t_{0}:t-1})$ (13)
-As described in Section B, within the denoising diffusion model, we need to train a neural network $\hat{\mathbf{x}}_{\theta_{3}}({}_{k}\mathbf{x}^{n}_{t},n,{}_{k}\mathbf{h}_{t_{0}:t-1})$ that predicts ${}_{k}\mathbf{x}^{0}_{t}$ from noisy vector ${}_{k}\mathbf{x}^{n}_{t}$, noise step $n$, and context embedding ${}_{k}\mathbf{h}_{t_{0}:t-1}$.
+As described in the previous section, within the denoising diffusion model, we need to train a neural network $\hat{\mathbf{x}}_{\theta_{3}}({}_{k}\mathbf{x}^{n}_{t},n,{}_{k}\mathbf{h}_{t_{0}:t-1})$ that predicts $_{k}\mathbf{x}^{0}_{t}$ from noisy vector $_{k}\mathbf{x}^{n}_{t}$, noise step $n$, and context embedding $_{k}\mathbf{h}_{t_{0}:t-1}$.
 We use a conditional UNet for this purpose.
-Within the UNet architecture, we employ a cross-attention conditioning mechanism (Rombach et al., 2022) to map the context embedding to the intermediate layers of the UNet.
-It allows us to predict ${}_{k}\mathbf{x}^{0}_{t}$ while paying attention to different parts of the context embedding.
-5.1 Training
-Input: Data: ${}_{k}\mathbf{x}^{0}_{t}\sim q({}_{k}\mathbf{x}^{0}_{t})$, Context: ${}_{k}{\mathbf{x}^{0}_{t_{0}:t-1}}$, ${}_{k}\mathbf{z}_{t_{0}:t-1}$, ${}_{k}\mathbf{s}$.
+Within the UNet architecture, we employ a cross-attention conditioning mechanism (Rombach_2022_CVPR) to map the context embedding to the intermediate layers of the UNet.
+It allows us to predict $_{k}\mathbf{x}^{0}_{t}$ while paying attention to the context embedding.
+Training
+Input: Data: $_{k}\mathbf{x}^{0}_{t}\sim q({}_{k}\mathbf{x}^{0}_{t})$, Context: $_{k}{\mathbf{x}^{0}_{t_{0}:t-1}}$, $_{k}\mathbf{z}_{t_{0}:t-1}$, $\mathbf{s}_{k}$.
 1 while not converged do
+ // Masking
+2   Randomly generate a binary sensor mask $\mathbf{M}_{k}$.
+3 for $\tau\in[t_{0},t-1]$ do
+4    $\mathbf{u}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$
+5    $_{k}\mathbf{y}^{0}_{\tau}={}_{k}\mathbf{x}^{0}_{\tau}\odot\mathbf{M}_{k}+({\mathbf{1}}-\mathbf{M}_{k})\odot\mathbf{u}$
+ // Compute context embedding
+[…]
+7 $_{k}\mathbf{g}_{t_{0}:t-1}\leftarrow\textsc{sinusoid}{}({}_{k}\mathbf{z}_{t_{0}:t-1})$
+[…]
  // Forward diffusion
@@
 10 $_{k}\mathbf{x}^{n}_{t}\leftarrow\mathcal{N}\left(\sqrt{\bar{\alpha}_{n}}\ {}_{k}\mathbf{x}^{0}_{t},(1-\bar{\alpha}_{n})\mathbf{I}\right)$.
-       // Compute context embedding
-4       ${}_{k}\mathbf{f}^{0}_{t_{0}:t-1}\leftarrow\textsc{conv}{}_{\theta_{1}}({}_{k}\mathbf{x}^{0}_{t_{0}:t-1}\oplus{}_{k}\mathbf{s})$
-5       ${}_{k}\mathbf{g}^{0}_{t_{0}:t-1}\leftarrow\textsc{sinusoid}{}({}_{k}\mathbf{z}_{t_{0}:t-1})$ ${}_{k}\mathbf{h}_{t_{0}:t-1}\leftarrow\textsc{linear}{}_{\theta_{2}}({}_{k}\mathbf{f}_{t_{0}:t-1}\oplus{}_{k}\mathbf{g}_{t_{0}:t-1})$
  // prediction
 11 $_{k}\hat{\mathbf{x}}^{0}_{t}\leftarrow\hat{\mathbf{x}}_{\theta_{3}}({}_{k}\mathbf{x}_{t}^{n},n,{}_{k}\mathbf{h}_{t_{0}:t-1})$
-7       Take gradient step on $\nabla_{\theta}||{}_{k}\mathbf{x}^{0}_{t}-{}_{k}\hat{\mathbf{x}}^{0}_{t}||^{2}_{2}$
-Algorithm 1 Training for patch $P_{k}$, prediction time step $t$
-To train Diff-Flood, we sample context windows (${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$) and the immediate next timestep (${}_{k}\mathbf{x}^{0}_{t}$) from the training time series data of each patch.
-We then optimize the parameters $\theta$ by minimizing the negative log-likelihood of the data ($-\log p_{\theta}({}_{k}\mathbf{x}^{0}_{t}|{}_{k}\mathbf{h}_{t_{0}:t-1}$).
-Through a similar derivation as shown in the Section B, the objective to minimize becomes:
-$\mathbb{E}_{{}_{k}\mathbf{x}^{0}_{t},n}||{}_{k}\mathbf{x}^{0}_{t}-\hat{\mathbf{x}}_{\theta}({}_{k}\mathbf{x}_{t}^{n},n,{}_{k}\mathbf{h}_{t_{0}:t-1}||^{2}_{2}$ (15)
-Here, the neural network $\hat{x}_{\theta}$ is now also conditioned on the context embedding ${}_{k}\mathbf{h}_{t_{0}:t-1}$.
-Input: Noise: ${}_{k}\mathbf{x}^{N}_{t}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$, Context: ${}_{k}{\mathbf{x}^{0}_{t_{0}:t-1}}$, ${}_{k}\mathbf{z}_{t_{0}:t-1}$, ${}_{k}\mathbf{s}$.
+[…]
+Algorithm 1 Training step on a data point of $P_{k}$.
+Input: Noise: $_{k}\mathbf{x}^{N}_{t}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$, Context: $_{k}{\mathbf{y}^{0}_{t_{0}:t-1}}$, $_{k}\mathbf{z}_{t_{0}:t-1}$, $\mathbf{s}_{k}$, and $\mathbf{M}_{k}$.
 1 for $n\in{N,...,1}$ do
+2 $\mathbf{u}\leftarrow{\mathbf{0}}$
 3 if $n>1$ then
-3             $\mathbf{v}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$
-4      else
-5             $\mathbf{v}\leftarrow{\mathbf{0}}$
+4    $\mathbf{u}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$
 5  Compute context embedding $_{k}\mathbf{h}_{t_{0}:t-1}$.
@@
 7 $_{k}\mathbf{x}^{n-1}_{t}\leftarrow\frac{\sqrt{\alpha_{n}}(1-\bar{\alpha}_{n-1})}{1-\bar{\alpha_{n}}}{}_{k}\mathbf{x}^{n}_{t}+\frac{\sqrt{\bar{\alpha}_{n-1}}\beta_{n}}{1-\bar{\alpha_{n}}}{}_{k}\hat{\mathbf{x}}^{0}_{t}$
-9       $\hskip 160.00024pt\ignorespaces+\sqrt{\Sigma_{\theta}}\mathbf{v}$
+8 $\hskip 160.00024pt\ignorespaces+\sqrt{\Sigma_{\theta}}\mathbf{u}$
 return $_{k}\mathbf{x}^{0}_{t}$
 Algorithm 2 Sampling $_{k}\mathbf{x}^{0}_{t}$
-5.2 Inference
-During inference, given a context ${}_{k}\mathbf{x}^{0}_{t_{0}:t-1},{}_{k}\mathbf{z}_{t_{0}:t-1},{}_{k}\mathbf{s}$, we first compute the context embedding ${}_{k}\mathbf{h}_{t_{0}:t-1}$.
+To train Diff-Sparse, we sample the following from the training time-series data of each patch: ($i$) inundation history in the context window ($_{k}\mathbf{x}^{0}_{t_{0}:t-1}$) and temporal co-variates ($_{k}\mathbf{z}_{t_{0},t-1}$), (iiii) inundation values in the immediate next timestep ($_{k}\mathbf{x}^{0}_{t}$).
+Diff-Sparse is trained to forecast the latter.
+We apply Algorithm 1 on each sample.
+First, we generate a random binary sensor mask $\mathbf{M}_{k}$.
+This mask is applied on the context inundation history ($_{k}\mathbf{x}^{0}_{t_{0}:t-1}$) to generate a sparse inundation history $_{k}\mathbf{y}^{0}_{t_{0}:t-1}$ (line 1-1, $\odot$ denotes element-wise multiplication).
+Here, we retain the inundation values in cells where sensors are placed (according to the sensor mask $\mathbf{M}_{k}$).
+For cells that do not have sensors, their inundation values are replaced with standard gaussian noise.
+This serves as a signal to the model to ignore these values.
+As we train Diff-Sparse with different random sensor masks, it is able to make forecasts based on different sensor placement configurations during inference.
+We use the sparse inundation history, temporal co-variates and the sensor mask to calculate the context embedding $_{k}\mathbf{h}_{t_{0}:t-1}$ (lines 1-1).
+We then optimize the model parameters $(\theta_{1},\theta_{2},\theta_{3})$ by minimizing the negative log-likelihood of the data conditioned on the context embedding ($\log p_{\theta}({}_{k}\mathbf{x}^{0}_{t}|{}_{k}\mathbf{h}_{t_{0}:t-1}$).
+Through a similar derivation as shown in the previous section, the objective to minimize becomes:
+$\mathbb{E}_{{}_{k}\mathbf{x}^{0}_{t},n}||{}_{k}\mathbf{x}^{0}_{t}-\hat{\mathbf{x}}_{\theta_{3}}({}_{k}\mathbf{x}_{t}^{n},n,{}_{k}\mathbf{h}_{t_{0}:t-1}||^{2}_{2}$ (14)
+Here, the neural network $\hat{x}_{\theta_{3}}$ is now also conditioned on the context embedding $_{k}\mathbf{h}_{t_{0}:t-1}$.
+Inference
+During inference, given a context ($_{k}\mathbf{y}^{0}_{t_{0}:t-1}$, $_{k}\mathbf{z}_{t_{0}:t-1}$, $\mathbf{s}_{k}$, $\mathbf{M}_{k}$), we first compute the context embedding $_{k}\mathbf{h}_{t_{0}:t-1}$.
 We then apply Algorithm 2 to obtain a sample $_{k}\mathbf{x}^{0}_{t}$ of the next time step.
@@
 We repeat the process multiple times to get multiple predictions.
-6 Experiment
-We conduct our experiments in two parts.
-First, we evaluate Diff-Flood model on the TideWatch dataset (see Section 6.1) against five competitive baseline models, taking forecasting inundation across the Virginian Eastern Shore region as our case study.
-Second, we conduct an ablation study of our model, underscoring the importance of the components of our proposed architecture.
-6.1 Dataset
-To train Diff-Flood, we downloaded inundation data of the Eastern Shore, Virginia from Tidewatch Maps (Loftis et al., 2019; Loftis, 2022); a street-level inundation mapping tool developed by the Virginia Institute of Marine Science (VIMS).
-It utilizes the open-source hydro-dynamic model, SCHISM (Zhang et al., 2016) as the engine, and provides 36 hour inundation forecast maps with a 12 hour update frequency.
+Experiment
+In our experiments, we first evaluate Diff-Sparse on the TideWatch dataset against six competitive baseline models; taking forecasting inundation within the Virginia Eastern Shore region as our case study.
+Then, we conduct an ablation study to demonstrate the importance of the different context information utilized in Diff-Sparse.
+Dataset
+To train Diff-Sparse, we downloaded inundation data of the Eastern Shore, Virginia from Tidewatch Maps (Loftis2019tidewatch; Loftis2022tidewatch); a street-level inundation mapping tool developed by the Virginia Institute of Marine Science (VIMS).
+It utilizes the open-source hydro-dynamic model, SCHISM (zhang2016seamless) as the engine, and provides 36 hour inundation forecast maps with a 12 hour update frequency.
 We used this dataset due to its high spatial resolution, coverage of areas affected by coastal inundation, and lack of other publicly available inundation datasets.
-We downloaded 89 days (March 28, 2024 to June 24, 2024) of hourly inundation data of the area.
+Moreover, using it as training data allows Diff-Sparse to learn to make forecasts from a physics-based hydro-dynamic model.
+The dataset will be shared upon request.
+We downloaded 89 days (March 28, 2024 to June 24, 2024) of hourly inundation data of our study area.
 The data comes in a raster geo-database format with a pixel resolution of $10m\times 10m$ and provides inundation level at each pixel in feet.
 We resampled the data to a pixel resolution of $30m\times 30m$.
-We then consider different square-shaped patches of size: $16\times 16$, $64\times 64$, $80\times 80$, and $96\times 96$.
-Each patch $P_{k}$ has a time series of inundation levels denoted by ${}_{k}\mathbf{x}^{0}_{t}\in\mathbb{R}^{D\times D}$, where $D$ is the length of one side of the patch.
-We aim to train Diff-Flood to forecast inundation levels on multiple patches.
-We have used 81 days (1944 timesteps, each of length 1-hour) of data for training, 1 day (24 timesteps) of data for validation, and 7 days (168 timesteps) of data for testing.
-Note that, when we utilize $K$ number of patches, the total number of both training and testing timesteps are multiplied by $K$.
-In our forecasting task, we use context window of $12$ hours, and then predict the inundation levels in the subsequent $12$ hours (prediction window).
-Each sample datapoint therefore contains 24 consecutive timesteps, first $12$ hours for context and the last $12$ hours for prediction.
-We normalize each datapoint separately as follows: first we calculate the mean and standard deviation of the inundation values in the context window, then we standardize the inundation values in both the context and prediction window using the calculated mean and standard deviation.
-In addition to the inundation data, we used digital elevation data (OpenTopography, 2021) for the same area.
-We resampled the elevation data to the same resolution of $30m\times 30m$ and then aligned it with our inundation data.
-6.2 Evaluation Metric
-We use two metrics to evaluate forecast quality with respect to the observations.
+We then consider different square-shaped patches.
+Each patch $P_{k}$ has a time series of inundation levels denoted by $_{k}\mathbf{x}^{0}_{t}\in\mathbb{R}^{D\times D}$, where $D\in\{16,64,80,96\}$ is the side length of the patch.
+We also used digital elevation data (DigitalElevation2021) for the same area.
+We resampled the elevation data to the same resolution and then aligned it with our inundation data.
+It serves as spatial context (i.e., elevation of patch $P_{k}$ denoted as $\mathbf{s}_{k}$) within Diff-Sparse.
+We have used 81 days of data for training, 1 day of data for validation, and 7 days of data for testing.
+Training data points are sampled by following the process described in Section ‘Diff-Sparse Method (Training)’.
+We use a context window of $12$ hours.
+During testing, we predict the inundation levels in the subsequent $12$ hours.
+Inundation values are standardized based on the mean and standard deviation of the inundation values in the training data.
+Elevation values are standardized separately based on the mean and standard deviation of the elevation values in the study area.
+Patch
+$D^{2}\times K$
+Metric DiffSTG DCRNN DeepVAR LSTNet TimeGrad BayesNF Diff-Sparse
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 1: Performance Comparison of Diff-Sparse with baseline models at 95% sparsity level (i.e., only $5\%$ of cells in a patch have sensors).
+All metrics are reported up to one standard deviation after conducting ten experiments with different random seeds.
+For each configuration, the best performance is bold-highlighted and the second-best performance is underlined.
+For Diff-Sparse, percentage improvement in performance with its closest baseline competitor is shown.
+State-of-the-art spatiotemporal forecasting methods (e.g., DiffSTG, DCRNN) could not handle training for large patch configurations; these are marked as ‘Failed’.
+Since, DCRNN and LSTNet are point forecasting methods, their NACRPS metrics are marked as ‘N/A’.
+Evaluation Metric
+We use two metrics to evaluate forecast quality.
 The first metric is Normalized Root Mean Squared Error (NRMSE).
-Let, $\mathbf{x}$ be the set of all observations over all cells for all timesteps.
-$x_{ij,t}\in\mathbf{x}$ be the observed value of cell $(i,j)$ at time $t$ and let, $\tilde{x}_{ij,t}$ denote the corresponding average forecast value.
-Then, NRMSE is defined as follows:
-$\text{NRMSE}=\frac{1}{\max_{ij,t}(\mathbf{x})-\min_{ij,t}(\mathbf{x})}\sqrt{\frac{\sum_{ij,t}{(x_{ij,t}-\tilde{x}_{ij,t})^{2}}}{\sum_{ij,t}1}}$ (16)
-In addition to NRMSE, we use an extended version of the Continuous Ranked Probability Score (CRPS) Winkler et al.
-(1996) to quantify forecast uncertainty.
-The CRPS metric is calculated for each cell and each timestep individually.
-Let’s assume that we have $M$ forecasts ${\hat{x}_{ij,t}^{1},\hat{x}_{ij,t}^{2},...,\hat{x}_{ij,t}^{M}}$ for observation $x_{ij,t}$.
-First, the empirical cumulative distribution function (CDF) at a point $z$, $\hat{F}_{ij,t}^{M}(z)$ from these forecast values is defined as
-$\hat{F}_{ij,t}^{M}(z)=\frac{1}{M}\sum_{m=1}^{N}{\mathbb{I}(\hat{x}_{ij,t}^{n}\leq z)}$ (17)
-where $\mathbb{I}(y)$ is the indicator function that yields $1$ if condition $y$ holds, 00 otherwise.
-From this, CRPS between the empirical CDF $\hat{F}_{ij,t}^{M}$ and the observation $x_{ij,t}$ is calculated as
+Let $\mathbf{x}$ be the set of all observations over all cells for all timesteps.
+Let $x_{ij,t}\in\mathbf{x}$ denote the observed inundation value of cell $(i,j)$ at time $t$ and let, $\tilde{x}_{ij,t}$ denote the corresponding average forecast value.
+Then:
+$\text{NRMSE}=\tfrac{1}{\max_{ij,t}(\mathbf{x})-\min_{ij,t}(\mathbf{x})}\sqrt{\tfrac{\sum_{ij,t}{(x_{ij,t}-\tilde{x}_{ij,t})^{2}}}{\sum_{ij,t}1}}$ (15)
 […]
-Since CRPS is defined for individual timestep and individual cells, we extend this metric to Normalized Average CRPS (NACRPS), defined as
-$\text{NACRPS}(\hat{F}^{M},\mathbf{x})=\frac{\sum_{ij,t}{\text{CRPS}(\hat{F}_{ij,t}^{M},x_{ij,t})}}{\sum_{ij,t}{|x_{ij,t}|}}$ (19)
+$\text{NACRPS}(\hat{F}^{M},\mathbf{x})=\tfrac{\sum_{ij,t}{\text{CRPS}(\hat{F}_{ij,t}^{M},x_{ij,t})}}{\sum_{ij,t}{|x_{ij,t}|}}$ (16)
 Here, $|x_{ij,t}|$ denotes the absolute value of $x_{ij,t}$.
-6.3 Baseline Comparison
-Patch
-$D\times D\times K$
-Metric DiffSTG DCRNN DeepVAR LSTNet TimeGrad Diff-Flood
-$16\times 16\times 2$ NACRPS 1.2523$\pm$0.0445 N/A 1.4782$\pm$0.0332 N/A 1.2319$\pm$0.0672 0.9603$\pm$0.2119 ($\uparrow$22.05%)
-NRMSE 0.1278$\pm$0.0022 0.1302$\pm$0.0125 0.1742$\pm$0.0034 0.1439$\pm$0.0003 0.1601$\pm$0.0071 0.1162$\pm$0.0189 ($\uparrow$9.02%)
-$16\times 16\times 5$ NACRPS 1.3166$\pm$0.0164 N/A 1.4151$\pm$0.0293 N/A 1.2558$\pm$0.0707 0.834$\pm$0.1781 ($\uparrow$33.59%)
-NRMSE 0.1373$\pm$0.0015 0.1166$\pm$0.0096 0.1734$\pm$0.0023 0.1501$\pm$0.0001 0.17$\pm$0.0075 0.1095$\pm$0.0165 ($\uparrow$6.13%)
-$64\times 64\times 10$ NACRPS Failed N/A 1.0228$\pm$0.0895 N/A 0.5522$\pm$0.0912 0.2007$\pm$0.0201 ($\uparrow$63.66%)
-NRMSE 0.0801$\pm$0.0065 0.3437$\pm$0.035 0.1765$\pm$0.0002 0.1818$\pm$0.0182 0.0639$\pm$0.0079 ($\uparrow$20.21%)
-$64\times 64\times 20$ NACRPS Failed Failed 1.1451$\pm$0.0792 N/A 0.545$\pm$0.0181 0.227$\pm$0.0249 ($\uparrow$58.34%)
-NRMSE 0.3967$\pm$0.0419 0.1941$\pm$0.0006 0.1908$\pm$0.0029 0.0794$\pm$0.008 ($\uparrow$58.39%)
-$80\times 80\times 20$ NACRPS Failed Failed 1.1494$\pm$0.1409 N/A 0.5251$\pm$0.0105 0.2091$\pm$0.0216 ($\uparrow$60.18%)
-NRMSE 0.2761$\pm$0.0399 0.131$\pm$0.0003 0.1251$\pm$0.0021 0.0508$\pm$0.0034 ($\uparrow$59.39%)
-$96\times 96\times 20$ NACRPS Failed Failed 1.3578$\pm$0.0372 N/A 0.5788$\pm$0.016 0.2519$\pm$0.0241 ($\uparrow$56.48%)
-NRMSE 0.1661$\pm$0.0057 0.0748$\pm$0.0001 0.0718$\pm$0.001 0.0334$\pm$0.0044 ($\uparrow$53.44%)
-Table 1: Performance Comparison of Diff-Flood with baseline models.
-All metrics are reported up to one standard deviation after conducting five experiments with different random seeds.
-For each configuration, the best performance is bold-highlighted and the second-best performance is underlined.
-For Diff-Flood, percentage improvement in performance with its closest baseline competitor is shown.
-State-of-the-art spatiotemporal forecasting methods (e.g., DiffSTG, DCRNN) could not handle training for large patch configurations, these are marked as ‘Failed’.
-Since, DCRNN and LSTNet are point forecasting methods, their NACRPS metric are marked as ‘N/A’.
-In this section, we describe the baseline models compared with our model.
-All experiments were conducted on an HPC cluster.
-To achieve scalability, experiments with different configurations were conducted with the SLURM scheduler.
-Each job used up to $256$GB CPU RAM, $8$ CPU cores and could request for one GPU with maximum $40$ GB GPU RAM.
+Baseline Comparison
+We now describe our baseline methods.
+All our experiments were conducted on an HPC cluster using the SLURM scheduler.
+Each job used up to 256256GB CPU RAM, $8$ CPU cores and one GPU with maximum $40$ GB GPU RAM.
+Our source code is provided with the supplementary materials.
   1. 1.
-    DiffSTG (Wen et al., 2023): The State-of-the-art (SOTA) diffusion-based model for spatiotemporal forecasting.
+    DiffSTG (wen2023diffstg): The State-of-the-art (SOTA) diffusion-based model for spatiotemporal forecasting.
   2. 2.
-    TimeGrad (Rasul et al., 2021) The SOTA diffusion-based model designed for Multivariate time-series forecasting.
+    TimeGrad (rasul2021autoregressive) The SOTA diffusion-based model designed for multivariate time-series forecasting.
   3. 3.
-    DeepVAR (Salinas et al., 2019): Combines an RNN-based architecture with a Gaussian copula process output model with a low-rank covariance structure.
+    DeepVAR (salinas2019high): Combines an RNN-based architecture with a Gaussian copula process output model with a low-rank covariance structure.
   4. 4.
-    LSTNet (Lai et al., 2018): The Long-and Short-term Time-series network (LSTNet) captures local dependency pattern among variables and long-term temporal correlation among time-series trends using CNN and RNN.
+    LSTNet (lai2018modeling): Captures local dependency pattern among variables and long-term temporal correlation among time-series trends using CNN and RNN.
   5. 5.
-    DCRNN (Li et al., 2017): A spatiotemporal forecasting framework that combines Diffusion Convolution and Sequence-to-sequence architecture to capture temporal and spatial dependencies.
-Each model was trained on six different patch settings as shown in Table 1.
-All hyper-parameters of our model are provided in the supplementary material.
-It is to be noted that the number of variables increases quadratically with patch size and linearly with the number of patches.
-Since some models do not scale well [^1] It is worth mentioning we also attempted to use $D^{3}$ (Li et al., 2022) as a baseline.
-However, we it scaled poorly; causing memory error even for our lowest patch configuration.
-with increments in the number of variables, we limit $D=96$ for this set of experiments.
-Across all experiments, we set the context and prediction length to 12.
+    DCRNN (li2017diffusion): A spatiotemporal forecasting framework that combines diffusion convolution and sequence-to-sequence architecture to capture temporal and spatial dependencies.
+  6. 6.
+    BayesNF (saad2024scalable): SOTA DL method based on a Bayesian Neural Network that maps a multivariate space-time coordinate to a real-valued field.
+We used six different patch settings in our experiments as shown in Table  1 (visualization of patches provided in supplementary materials).
+The experiments are performed at a sparsity level of $95\%$; meaning when forecasting for a patch, only $5\%$ of the cells of that patch will have sensors and inundation values in the context window.
+During training, we generate random sensor masks ($\mathbf{M}_{k}$) that have $\sim 5\%$ of the cell values to be $1$.
+For testing, we generate $10$ different sensor masks at random with same ($95\%$) sparsity level for each patch setting.
+These masks are applied on test data points (from Tidewatch) in a round-robin manner; i.e., mask 1 applied on data point 1, mask 2 applied on data point 2, …, mask 10 applied on datapoint 10, mask 1 applied on datapoint 11, and so on.
+This ensures that all forecasting methods are being tested with same datapoints.
+By masking the test datapoints and then making forecasts based on them, we simulate the real-world scenario of forecasting based on sparse sensor observations.
+It should to be noted that the number of variables increases quadratically with patch size and linearly with the number of patches.
+We found that some models do not scale well [^1] We considered $D^{3}$VAE (li2022generative) as a baseline.
+It scaled poorly; causing memory error for our lowest patch configuration.
+with increments in the number of variables.
+Across all experiments, we set the context and prediction length to 12 (experiment results with varying prediction length are provided in the supplementary materials).
 Eight scenarios were sampled to understand the quality of probabilistic forecasting.
-Table 1 summarizes the performance of our model with respect to the baseline models.
-Below, we outline some key observations from the results.
-Figure 3: Scaling in Model Parameters (Log scale in Y axis) with increasing patch configuration.
-It has to be noted that model parameter is not affected by $K$.
-DiffSTG and DCRNN are not shown here since they fail to execute for large patch setting.
-First, in terms of NACRPS, our model consistently outperforms the baselines, underscoring its superiority in probabilistic forecasting.
-For small patch size ($D=16$), Diff-Flood has a performance gain of between $22.05\%$ and $33.59\%$, compared to its closest baseline.
-For larger patch sizes ($D\geq 64$), the improvement in NACRPS is close to $60\%$, underscoring the performance scalability of our model.
-Second, Diff-Flood outperforms other baselines in terms of NRMSE too.
-Although TimeGrad was constantly the closest competitor baseline in terms of NACRPS, this is not the case here.
-For the smallest patch setting, DiffSTG was the best baseline in terms of NRMSE, which was outperformed by Diff-Flood by $9.02\%$.
-For the next two patch settings, DCRNN was the best baseline.
-In these cases, Diff-Flood outperformed DCRNN by $6.13\%$ and $20.21\%$, respectively.
-For the settings with $K=20$, DiffSTG and DCRNN both required extensive GPU memory due to poor scalability and they failed to train.
-In these settings, Diff-Flood outperformed TimeGrad with a margin of around $50-60\%$.
-These experiments highlight that existing spatiotemporal forecasting methods (i.e., DiffSTG and DCRNN) have moderate performance but poor computation scalability, making them infeasible to deploy in large spatiotemporal settings.
-On the other hand, temporal forecasting methods do not have good performance scalability, making them inadequate for large-scale spatiotemporal forecasting tasks.
-Diff-Flood has both computation and performance scalability, making it suitable for large-scale spatiotemporal forecasting tasks.
-Finally, the number of parameters of the baseline models grows almost exponentially with increasing patch sizes ($D$), as observed from Figure 3.
-To investigate the reason, we analyzed the implementation of TimeGrad as an example and found that it uses a fully connected layer proportional to the number of variables.
-On the other hand, our model utilizes a UNet architecture that creates a low-dimensional latent representation of the patches.
-It then performs expensive operations (e.g., feed-forward, cross-attention) on the lower-dimensional latent space, allowing it to scale quite well with the number of variables.
-6.4 Ablation Study
-Figure 4: Bar-plots showing performance comparison of the four configurations of Diff-Flood on our test dataset of configuration $64\times 64\times 10$, in terms of two metrics.
-The height of the bars represent mean of the metric and the vertical black lines represent standard deviation.
-Lower metric value is better.
-Our proposed architecture of Diff-Flood is able to utilize elevation (ELEV) and temporal covariates (COV) individually, resulting in improved performance in terms of both of our two metrics.
-Best performance is achieved when both are used.
-In this section, we perform an ablation study of the Diff-Flood method to investigate the effects of utilizing elevation matrix and temporal covariates, in addition to the inundation history.
-Specifically, given inundation history (INUN: ${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$), elevation (ELEV: ${}_{k}\mathbf{s}$), and temporal covariates (COV: $\mathbf{z}_{t_{0}:t-1}$), we define four configurations based on how the context embedding is computed:
-  1. 1.
-    INUN: Only the historical inundation (${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$) used.
-  2. 2.
-    INUN, ELEV: We use historical inundation data ${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$ and the elevation matrix ${}_{k}\mathbf{s}$.
-  3. 3.
-    INUN, COV: We use historical inundation data ${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$ and the temporal covariates $\mathbf{z}_{t_{0}:t-1}$.
-  4. 4.
-    INUN, ELEV, COV: We include all three components.
-We ran five experiment runs for each configuration, where we trained and tested Diff-Flood on the Tidewatch dataset with $K=10$ and $D=64$.
-Figure 4 shows the performance of the four configurations in-terms of NACRPS and NRMSE.
-We observe that, configuration (INUN, COV) and (INUN, ELEV) are both better than the configuration (INUN) in terms of the two metrics.
-This implies that Diff-Flood is able to utilize both elevation and temporal covariates as context effectively.
-We also observe that, including the temporal covariates results in a greater improvement ($43\%,41\%$ improvement over INUN in terms of NACRPS and NRMSE, respectively) than including spatial context ($8\%,4\%$).
-However, best performance is achieved when both are utilized ($48\%,47\%$ improvement over INUN).
-7 Discussion and Conclusion
-Once Diff-Flood is trained, we can use it to answer inundation-scenario related queries relevant to policymakers. Example queries are as follows:
-Figure 5: (left) A school area in the Eastern Shore of Virginia.
-The red cells and four orange patches overlap with the school area.
-A possible query is: what is the probability that the flooding level in the school area will be above $d$ units within next $T$ hours?
-(right) A route coming out from the school.
-The red cells and the eight orange patches overlap with the route.
-A possible query is: what is the probability that the route coming out of the school will not be flooded in the next $T$ hours?
-  • Query 1: What is the probability that the flooding level in an area $A$ will be above $d$ units within next $T$ hours? (Figure 5, left)
-  • Query 2: Given a route, what is the probability that the route will not be flooded in the next $T$ hours? (Figure 5, right)
-Such queries are of special interest to city officials; answer to these queries will help to better prepare and respond to emergencies.
-The key idea for answering such queries is that, we have a sampling method in the form a trained Diff-Flood model and we can use it to sample inundation scenarios from the distribution of possible scenarios.
-We use the sampled scenarios to calculate our desired probability value.
-In the supplementary materials, we provide a detailed algorithm for answering such queries using a trained Diff-Flood model.
-Although Diff-Flood is developed to solve the task of coastal inundation forecasting, it is generalizable enough to be extended to other spatiotemporal forecasting tasks (e.g., occupancy detection, traffic forecasting).
-Diff-Flood supports conditioning on time-invariant spatial features, as well as, on temporal co-variates.
-In our future work, we plan to evaluate its performance in diverse spatiotemporal forecasting tasks.
-To conclude, we have presented a probabilistic spatiotemporal forecasting method, Diff-Flood that takes both spatial and temporal context into account during its training and forecasting phase.
-We have performed a case study on the coastal inundation forecasting task for the Eastern Shore of Virginia.
-Through our experiments, we show that Diff-Flood outperforms existing forecasting methods in predicting coastal inundation.
-Additionally, the architecture of Diff-Flood allows it to scale to large forecasting tasks at fine spatial resolution, compared to current methods.
-Appendix A Extended Related Work
-Hydrology Models Traditionally, inundation has been simulated through hydro-dynamical physics-based 1D/2D dual drainage models.
+Table 1 summarizes the performance of Diff-Sparse with respect to the baseline methods.
+Some key observations from the results are as follows:
+First, we observe that DiffSTG, DCRNN and BayesNF scale poorly to large patch configurations.
+DiffSTG was unable to process patch configurations where $D>16$.
+For DCRNN, $(D=64,K=10)$ was the highest patch configuration that was processed successfully, whereas BayesNF fails for $D>64$.
+In larger settings, we ran out of GPU memory for these methods.
+It demonstrates the relatively worse computational scalability of these three methods.
+Second, in small patch configurations, specifically $D=16\text{ and }K\in\{2,5\}$, Diff-Sparse performs second best.
+Here, BayesNF is the superior model.
+Diff-Sparse outperforms the other baselines in these configurations.
+Third, in moderate and large patch configurations ($D\in\{64,80,96\}$, $K\in\{10,20\}$), Diff-Sparse outperforms all six baseline methods in terms of both metrics.
+Across moderate configuration where $D=64$, which is the largest configuration that BayesNF scales to, Diff-Sparse outperforms BayesNF by a margin of $\sim 7$-$24\%$ and $\sim 18$-$29\%$ in terms of NACRPS and NRMSE, respectively.
+In larger configurations ($D\in\{80,96\},K=20$), Diff-Sparse achieves $59$-$61\%$ improvement over the second-best method (TimeGrad), in terms of the two performance metrics.
+This illustrates the superiority of Diff-Sparse over baseline methods in performance scalability.
+These experiment results highlight that existing spatiotemporal forecasting methods have either poor computational scalability or performance scalability, making it challenging to apply them to large-scale high-resolution spatiotemporal forecasting tasks.
+Diff-Sparse offers both computation and performance scalability, making it suitable for large-scale high-resolution spatiotemporal forecasting tasks under sparse observations.
+Figure 2: Bar-plots showing Diff-Sparse performance at varying sparsity levels in two patch configurations.
+Ten experiment runs were performed in each setting; mean and standard deviation of the performance metrics are shown.
+Ablation Study
+In this section, we perform ablation studies on Diff-Sparse to understand the importance of different context data.
+First, we investigate the predictive performance of Diff-Sparse at various sparsity levels.
+We trained and tested Diff-Sparse at four different sparsity levels ($0\%$, $50\%$$95\%$, and $100\%$) for the patch configurations $(64^{2},10)$ and $(64^{2},20)$.
+Sparsity level of $100\%$ means no sensor data is available.
+Figure 2 shows the performance of Diff-Sparse in these settings, in terms of our two performance metrics.
+We observe that in both patch configurations, Diff-Sparse has the best performance at $0\%$ sparsity level in terms of both metrics.
+As the sparsity level increases, performance of Diff-Sparse degrades; the worst performance observed at $100\%$ sparsity level.
+It implies that, having (more) sensors, i.e., observed inundation values in the context window, helps Diff-Sparse to make better predictions.
+Figure 3: Ablation study at varying sparsity levels for two patch configurations. Ten experiment runs were performed in each setting; mean value of the performance metrics are shown using bar-plots.
+Next, we investigate the utility of elevation data and temporal co-variates as additional context.
+Given sparse inundation history (INUN: $_{k}\mathbf{y}^{0}_{t_{0}:t-1}$), elevation (ELEV: $\mathbf{s}_{k}$), and temporal co-variates (COV: $\mathbf{z}_{t_{0}:t-1}$), we examine four configurations of Diff-Sparse based on how the context embedding is computed: (1) INUN: only sparse inundation values used, (2) INUN, ELEV: sparse inundation values and elevation data used, (3) INUN, COV: sparse inundation values and temporal co-variates used, (4) INUN, ELEV, COV: all three components used.
+Figure 3 shows the performance of these four settings at three different sparsity levels ($0\%$, $50\%$, and $95\%$) for the patch configurations $(64^{2},10)$ and $(64^{2},20)$.
+We see that in all sparsity levels, both (INUN, COV) and (INUN, ELEV) performs better than INUN; meaning both elevation data and temporal co-variates are individually helpful in making better forecasts.
+At low sparsity levels ($0\%$, $50\%$), the setting (INUN, COV) performs better than (INUN, ELEV), indicating the higher utility of temporal co-variates than elevation at low sparsity levels.
+However, at high sparsity level ($95\%$), we see the opposite; (INUN, ELEV) performs better than (INUN, COV).
+It means at high sparsity level, elevation data becomes more useful compared to temporal co-variates.
+In all sparsity levels, using all three components (i.e., INUN, COV, ELEV) yields the best result.
+Conclusion
+In this paper, we have addressed the problem of high resolution probabilistic coastal inundation forecasting from sparse sensor observations.
+To solve it, we first formulated it as a problem of learning a conditional distribution function.
+We then presented Diff-Sparse, a scalable spatiotemporal forecasting method that utilizes sparse sensor observations, elevation data, and temporal covariates to make probabilistic forecasts.
+Diff-Sparse is trained using data generated by a physics-based hydro-dynamical model.
+A novel masking strategy is employed to tackle the challenge of making forecasts based on sparse sensor observations during inference.
+Moreover, due to our robust training strategy, Diff-Sparse can make forecasts based on different sensor placement configurations without requiring retraining from scratch.
+Through our experiments, we demonstrated that Diff-Sparse outperforms existing forecasting methods in terms of computational and performance scalability.
+Our extensive ablation study reveals that all three spatial and temporal context data: elevation data, temporal covariates, and sensor observations (even if sparse), help Diff-Sparse make better forecasts.
+Acknowledgments
+This work was partially supported by the following grants: (i) NSF Grants CCF-1918656, OAC-1916805, RISE-2053013; (ii) AI Institute: Agricultural AI for Transforming Workforce and Decision Support (AgAID) award No.
+2021-67021-35344; (iii) U.S.
+Department of Energy, through the Office of Advanced Scientific Computing Research’s “Data-Driven Decision Control for Complex Systems (DnC2S)” project.
+Pacific Northwest National Laboratory is operated by Battelle Memorial Institute for the U.S.
+Department of Energy under Contract No.
+DE-AC05-76RL01830.
+Appendix A Related Work
+Hydrology Models: Traditionally, inundation has been simulated through hydro-dynamical physics-based 1D/2D dual drainage models.
 By simulating water flow in water bodies and coupling them with 1D or 2D hydrodynamic models, these models can predict inundation in surfaces.
-For example, the 2D model SOBEK (Deltares, 2014) consists of a 1D network with information about the initial water volume of rivers and 2D rectangular computational cells.
+For example, the 2D model SOBEK (deltares2014sobek) consists of a 1D network with information about the initial water volume of rivers and 2D rectangular computational cells.
 By coupling these two parts, the model uses the Saint-Venant flow equation and 2D shallow water equation to calculate discharge between cells and determine the water level at each cell of the Digital Elevation Model (DEM).
-TELEMAC-2D (Vu et al., 2015) also solves 2D shallow water equation to simulate water flow.
+TELEMAC-2D (vu2015two) also solves 2D shallow water equation to simulate water flow.
 However, its primary purpose is to understand flow through water bodies.
 Therefore, this makes this model more suitable for coastal regions, contrary to SOBEK which is more suitable for urban areas.
-Polymorphic models like SCHISM (Zhang et al., 2016) can capture water flow at very high resolution, down to even a few meters due to their hybrid grinding system and localized coordinate system.
-Despite these models’ ability to model inundation accurately at fine resolution, the computation time associated with them makes them infeasible for real-time forecasting.
+Polymorphic models like SCHISM (zhang2016seamless) can capture water flow at very high resolution, down to even a few meters due to their hybrid grinding system and localized coordinate system.
+Despite these models’ ability to model inundation accurately at fine resolution, the computation time associated with them makes them infeasible for real-time forecasting  (guo2021urban).
 Regression-based Models Early approaches in finding surrogates for hydrology models involved applying various regression-based methods.
-(Chang et al., 2010) first applied K-means clustering to pre-process data to find inundation control points.
+Chang et al.
+(chang2010clustering) first applied K-means clustering to pre-process data to find inundation control points.
 The authors then applied linear regression or neural networks to predict inundation for grids depending on whether they are linearly correlated with the control points or not.
-(Bermúdez et al., 2019) explored non-parametric regression approach based on least square support vector machine (SVM) as surrogate for 2D hydrodynamic models.
-(Lee and Kim, 2021) created an offline rainfall-runoff-inundation database by simulating hydrology models over multiple scenarios.
+Bermúdez et al.
+(bermudez2019rapid) explored non-parametric regression approach based on least square support vector machine (SVM) as surrogate for 2D hydrodynamic models.
+Lee et al.
+(lee2021scenario) created an offline rainfall-runoff-inundation database by simulating hydrology models over multiple scenarios.
 Then, they trained a logistic regression-based approach based on the database to estimate the parameters to predict the risk of flooding for each region.
 Their goal was then to use the trained regressor to generate near real-time prediction of inundation probability given the runoff to the grid from simulated and realistic rainfall scenario.
-Deep Learning (DL) Models Many DL models have been explored as possible surrogates for Physics Based Models (PBM).
-One research work (Pan et al., 2011) applied a combined approach consisting of principal component analysis (PCA) and a feed-forward neural network.
+Deep Learning: DL methods are a popular choice in surrogating hydrodynamic models.
+Pan et al.
+(pan2011hybrid) applied a combined approach consisting of principal component analysis (PCA) and a feed-forward neural network.
 Their approach involved inundation prediction at various locations in the next timestep by looking at a fixed window of rainfall history of $N$ different rainfall gauges.
 First, they applied PCA on these various rainfall gauge histories to reduce the dimensionality before using them as input to the feed-forward neural network.
-Since such neural networks are expected to fail where data availability is limited, another work (Xie et al., 2021) explored the potential of a hybrid feed-forward neural network-based approach where they used two block-based neural networks along with the traditional point-based neural network and found superiority of the hybrid approach over other in the data-sparse region.
-Consequently, convolutional neural network (CNN) was explored (Kabir et al., 2020) as a potential surrogate.
+Since such neural networks are expected to fail where data availability is limited, Xie et al.
+(xie2021artificial) explored the potential of a hybrid feed-forward neural network-based approach where they used two block-based neural networks along with the traditional point-based neural network and found superiority of the hybrid approach over other in the data-sparse region.
+Consequently, convolutional neural network (CNN) was explored (kabir2020deep) as a potential surrogate.
 After their model was trained to predict the output simulated from a 2D hydrology model, they used it to simulate two real-event flood scenarios in a region of UK.
-Since these models are not able to capture long-term temporal dependencies, researchers have recently explored (Roy et al., 2023) seq2seq LSTM model as surrogates to hydrology models.
+Since these models are not able to capture long-term temporal dependencies, researchers have recently explored (roy2023application) seq2seq LSTM model as surrogates to hydrology models.
 Such model was also applied in forecasting for multi-step future horizons.
-Diffusion Model TimeGrad (Rasul et al., 2021) was the pioneering work that proposed the use of diffusion models for time-series forecasting tasks.
+Aside from this, regression-based methods (zahura2022predicting) and hybrid ML-based methods (zanchetta2022hybrid) have been attempted as possible surrogates.
+However, these models cannot inherently take spatial context into account.
+The DCRNN method proposed by (li2017diffusion) for traffic forecasting is the state-of-the-art DL method for spatiotemporal point forecasting, where bidirectional random walk is used to capture spatial dependency and Seq2Seq architecture is used to capture temporal dependency.
+Diffusion Model The field of generative modeling has been dominated by the diffusion model for quite some time.
+Although it had primarily been studied for image synthesis (ho2020denoising; dhariwal2021diffusion; austin2021structured), its fascinating ability to use flexible conditioning to guide the generation process has been utilized in other domains like video generation (yang2023diffusion), prompt-based image generation (ramesh2022hierarchical), text-to-speech generation (yang2023diffsound).
+TimeGrad (rasul2021autoregressive) was the pioneering work that proposed the use of diffusion models for time-series forecasting tasks.
 They used temporal co-variates and embedded historical data as conditions for the diffusion model to denoise the prediction for the next time step.
-CSDI Tashiro et al.
-(2021) does the same but instead of encoding with a Recurrent Neural Network (RNN) as done in TimeGrad, they leverage a transformer-based architecture.
-SSSDS4 Alcaraz and Strodthoff (2022) uses a S4 model Gu et al.
-(2021) to capture the encoding which is particularly useful in capturing long-term dependencies.
+CSDI (tashiro2021csdi) does the same but instead of encoding with a Recurrent Neural Network (RNN) as done in TimeGrad, they leverage a transformer-based architecture.
+SSSDS4 (alcaraz2022diffusion) uses an S4 model (gu2021efficiently) to capture the encoding which is particularly useful in capturing long-term dependencies.
 Also, instead of performing diffusion along both temporal and input channel dimensions, they perform diffusion only along the temporal dimension.
-Different from these, $D^{3}$ (Li et al., 2022) uses coupled diffusion process and bidirectional variational autoencoder to better forecast in longer horizon and improve interpretability.
-The State-of-the-Art Diffusion-based model DiffSTG (Wen et al., 2023) devised a modified architecture for the reverse process of diffusion model combining UNet and GNN to capture temporal and spatial dependencies, respectively.
+Other models like DPSD-CSPD (bilovs2023modeling), TDSTF (Chang2024) were later introduced.
+These models try to predict the noise added during the forward process of the diffusion model ($\epsilon$-parameterization).
+An alternative parameterization is to predict the ground-truth data ($x$-parameterization).
+Different from models mentioned above, $D^{3}$VAE (li2022generative) uses coupled diffusion process and bidirectional variational autoencoder to better forecast in longer horizon and improve interpretability.
+The State-of-the-Art Diffusion-based model DiffSTG (wen2023diffstg) devised a modified architecture for the reverse process of diffusion model combining UNet and GNN to capture temporal and spatial dependencies, respectively.
 However, both these methods scale poorly with the growth in number of variables.
+BayesNF (saad2024scalable) is the state-of-the-art DL spatiotemporal forecasting model, which is built using a Bayesian Neural Network.
+By assigning a prior distribution to the model parameters and adjusting the posterior based on observation, BayesNF essentially learns to map multivariate spatiotemporal points to a continuous real-valued field.
 Appendix B Denoising Diffusion Model
 In this section, we provide an overview of denoising diffusion models.
-[…]
+Let $\mathbf{x}^{0}\sim q_{\chi}(\mathbf{x}^{0})$ denote the multivariate training vector from some input space $\chi=\mathbb{R}^{D}$ and true distribution $q_{\chi}(\mathbf{x}^{0})$.
+In diffusion probabilistic models, we aim to approximate $q_{\chi}(\mathbf{x}^{0})$ by a probability density function $p_{\theta}(\mathbf{x}^{0})$ where $\theta$ denotes the set of trainable parameters.
+Diffusion probabilistic models (sohl2015deep; luo2022understanding) are a special class of Hierarchical Variational Autoencoders (kingma2016improved; sonderby2016ladder) with the form $p_{\theta}(\mathbf{x}^{0})=\int p_{\theta}(\mathbf{x}^{0:N})d\mathbf{x}^{1:N}$; here $\mathbf{x}^{1},\mathbf{x}^{2},...,\mathbf{x}^{N}$ are latent variables.
+Three key properties of diffusion models are: $(i)$ all of the latent variables are assumed to have the dimension $D$ as $\mathbf{x}^{0}$, $(ii)$ the approximate posterior $q(\mathbf{x}^{1:N}|\mathbf{x}^{0})$,
 $q(\mathbf{x}^{1:N}|\mathbf{x}^{0})=\prod_{n=1}^{N}q(\mathbf{x}^{n}|\mathbf{x}^{n-1})$
@@
 This is done by minimizing the negative log-likelihood via a variational bound:
-[…]
+$\log p_{\theta}(\mathbf{x}^{0})$ $=\log\int p_{\theta}(\mathbf{x}^{0:N})d\mathbf{x}^{1:N}$
 $=\log\mathbb{E}_{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}\frac{p_{\theta}(\mathbf{x}^{0:N})}{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}$
@@
 Here, we get (18) by applying Jensen’s inequality.
-We indirectly minimize the negative log-likelihood of the data, i.e., $-\log p_{\theta}(\mathbf{x}^{0})$, by minimizing the right hand side of (22) .
+We indirectly minimize the negative log-likelihood of the data, i.e., $\log p_{\theta}(\mathbf{x}^{0})$, by minimizing the right hand side of (19) .
 We can show that it is equal to:
 $\mathbb{E}_{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}\left[-\log p_{\theta}(\mathbf{x}^{N})+\sum_{n=1}^{N}\log\frac{q(\mathbf{x}^{n}|\mathbf{x}^{n-1})}{p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})}\right]$ (20)
-Ho et al.
-(2020) showed that the forward process has the following property: we can sample $\mathbf{x}^{n}$ using $\mathbf{x}^{0}$ at any noise step $n$ in closed form .
+(ho2020denoising) showed that the forward process has the following property: we can sample $\mathbf{x}^{n}$ using $\mathbf{x}^{0}$ at any noise step $n$ in closed form .
 Let, $\alpha_{n}=1-\beta_{n}$, and $\bar{\alpha}_{n}=\prod_{i=1}^{n}\alpha_{i}$.
@@
 $D_{KL}\left(q(\mathbf{x}^{n-1}|\mathbf{x}^{n},\mathbf{x}^{0})\ ||\ p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})\right)$
-$=\frac{1}{2\tilde{\beta}_{n}}||\tilde{\mu}_{n}(\mathbf{x}^{n},\mathbf{x}^{0})-\mu_{\theta}(\mathbf{x}^{n},n)||_{2}^{2}+C$ (29)
+[…]
 Here, $C$ is a constant independent of $\theta$, and we assume that
@@
 $D_{KL}\left(q(\mathbf{x}^{n-1}|\mathbf{x}^{n},\mathbf{x}^{0})\ ||\ p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})\right)$
-$=\frac{1}{2\tilde{\beta}_{n}}\frac{\bar{\alpha}_{n-1}\ \beta_{n}^{2}}{(1-\bar{\alpha}_{n})^{2}}||\mathbf{x}^{0}-\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)||_{2}^{2}$ (32)
+[…]
 Therefore, optimizing a diffusion model can be done by training a neural network to predict the original ground truth vector from an arbitrarily noisy version of it.
 To approximately minimize the summation term in the objective (22) across all noise steps $n$, we can minimize the expectation over all noise steps as follows:
-$\mathop{\mathrm{argmin}}_{\theta}\ \mathbb{E}_{n\sim U[2,N]}\ \mathbb{E}_{q(\mathbf{x}^{n}|\mathbf{x}^{0})}\frac{1}{2\tilde{\beta}_{n}}\frac{\bar{\alpha}_{n-1}\beta_{n}^{2}}{(1-\bar{\alpha}_{n})^{2}}||\mathbf{x}^{0}-\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)||_{2}^{2}$ (33)
+[…]
 This can be done by using stochastic samples over noise steps.
@@
 Then, we compute:
-$\mathbf{x}^{n-1}=\mu_{\theta}(\mathbf{x}^{n},n)+\sqrt{\Sigma_{\theta}}\ \mathbf{v}$ (34)
-Here, $\mathbf{v}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$ for $n\in[2,N]$ and $\mathbf{v}={\mathbf{0}}$ when $n=1$.
+$\mathbf{x}^{n-1}=\mu_{\theta}(\mathbf{x}^{n},n)+\sqrt{\Sigma_{\theta}}\ \mathbf{u}$ (31)
+Here, $\mathbf{u}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$ for $n\in[2,N]$ and $\mathbf{u}={\mathbf{0}}$ when $n=1$.
 $\mu_{\theta}(\mathbf{x}^{n},n)$ and $\Sigma_{\theta}$ are computed using (28) and (27), respectively.
-Appendix C Inundation Scenario Queries
-Once Diff-Flood is trained, we can use it to answer inundation scenario-based queries relevant to policymakers. In this Section, we discuss a few of such queries.
+Appendix C Evaluation Metric
+We use two metrics to evaluate forecast quality.
+The first metric is Normalized Root Mean Squared Error (NRMSE).
+Let $\mathbf{x}$ be the set of all observations over all cells for all timesteps.
+Let $x_{ij,t}\in\mathbf{x}$ denote the observed inundation value of cell $(i,j)$ at time $t$ and let, $\tilde{x}_{ij,t}$ denote the corresponding average forecast value.
+Then:
+$\text{NRMSE}=\frac{1}{\max_{ij,t}(\mathbf{x})-\min_{ij,t}(\mathbf{x})}\sqrt{\frac{\sum_{ij,t}{(x_{ij,t}-\tilde{x}_{ij,t})^{2}}}{\sum_{ij,t}1}}$ (32)
+In addition to NRMSE, we use an extended version of the Continuous Ranked Probability Score (CRPS) (winkler1996scoring) to quantify forecast uncertainty.
+The CRPS metric is calculated for each cell and each timestep individually.
+Let’s assume that we have $M$ forecasts $\hat{x}_{ij,t}^{1},\hat{x}_{ij,t}^{2},...,\hat{x}_{ij,t}^{M}$} for observation $x_{ij,t}$.
+First, the empirical cumulative distribution function (CDF) at a point $z$, $\hat{F}_{ij,t}^{M}(z)$ from these forecast values is defined as:
+$\hat{F}_{ij,t}^{M}(z)=\frac{1}{M}\sum_{m=1}^{N}{\mathbb{I}(\hat{x}_{ij,t}^{n}\leq z)}$ (33)
+where $\mathbb{I}(y)$ is the indicator function that yields $1$ if condition $y$ holds, 0 otherwise.
+From this, CRPS between the empirical CDF $\hat{F}_{ij,t}^{M}$ and the observation $x_{ij,t}$ is calculated as
+$\text{CRPS}(\hat{F}_{ij,t}^{M},x_{ij,t})=\int_{z}{\Big(\hat{F}_{ij,t}^{M}(z)-\mathbb{I}({x}_{ij,t}\leq z)\Big)^{2}dz}$ (34)
+Since CRPS is defined for individual timestep and individual cells, we extend this metric to Normalized Average CRPS (NACRPS), defined as
+$\text{NACRPS}(\hat{F}^{M},\mathbf{x})=\frac{\sum_{ij,t}{\text{CRPS}(\hat{F}_{ij,t}^{M},x_{ij,t})}}{\sum_{ij,t}{|x_{ij,t}|}}$ (35)
+Here, $|x_{ij,t}|$ denotes the absolute value of $x_{ij,t}$.
+Appendix D Additional Experiment Results
+Varying Prediction Length
+As Diff-Sparse forecasts the inundation levels in future timesteps auto-regressively, we can use it to forecast inundation levels for arbitrary number of future timesteps.
+However, this can result in error accumulation.
+To test this, we took the setting $(64^{2},10)$ and varied the prediction length with the values $1,4,12,20,28,36,44,52,60$.
+Figure 4 shows the values of the two performance metrics (NACRPS and NRMSE) for our test data, for each of these prediction lengths, where context length is set to 12 hours and sparsity level to $95\%$.
+We observe that Diff-Sparse performs best, in terms of both performance metrics, when prediction length is $1$.
+This is expected as Diff-Sparse is trained to make predictions for one future timestep.
+As the prediction length is increased, we observe that both NACRPS and NRMSE increases.
+Figure 4: Line-plots showing the performance of Diff-Sparse for different prediction lengths with test dataset configuration $(64^{2},10)$, context length of 12 hours, and sparsity level of $95\%$, in terms of two metrics.
+The dots represent the mean value of the metrics over ten experiment runs with different seeds; the vertical lines represent standard deviation.
+We observe that Diff-Sparse performs best when prediction length is $1$.
+This is expected as Diff-Sparse is trained to make predictions for one future timestep.
+When making predictions for multiple timesteps auto-regressively, we observe that both NACRPS and NRMSE increase with increasing number of prediction timesteps.
+Masking Strategy
+In Diff-Sparse, we have presented a novel masking strategy (Algorithm 1, lines 1-1).
+In this strategy, we retain the inundation values in cells where sensors are placed (according to the sensor mask).
+For cells that do not have sensors, their inundation values are replaced with standard gaussian noise.
+We refer to this masking strategy as ”Noise Masking”.
+To show the efficacy of this masking strategy, we consider a baseline masking strategy as follows: retain the inundation values in cells where sensors are placed; cells that do not have sensors, their inundation values are replaced with the value 0.
+We refer to this baseline masking strategy as ”Zero Masking”.
+Figure 5: Bar-plots showing the performance of Diff-Sparse (in terms of NACRPS and NRMSE) for two different masking strategies during training (Zero mask and Noise mask), with patch configuration $(64^{2},10)$ and $(64^{2},20)$ and sparsity level $95\%$.
+Context and prediction lengths are set to 12 hours.
+The height of the bars represent the mean value of the metrics over ten experiment runs with different seeds; the vertical black lines represent standard deviation.
+We observe that Diff-Sparse noise masking yields better result in terms of both performance metrics for the two patch configurations.
+Figure 5 shows a comparison of the two masking strategies in terms of the resulting predictive performance of Diff-Sparse.
+We observe that, for the patch configurations $(64^{2},10)$ and $(64^{2},20)$ at $95\%$ sparsity level, noise masking yields better predictive performance than zero masking.
+Our hypothesis is, by placing random gaussian noise, noise masking provides an additional signal to the model to ignore the inundation values where sensors are not available.
+Forecasting Time
+To understand the efficacy of Diff-Sparse in real-time forecasting, we recorded its forecasting time.
+With context length of 12 hours, prediction length of 36 hours, and number of scenarios to be predicted set to $8$; Diff-Sparse takes $\sim 7$ seconds to forecast for a patch of size $64\times 64$, on a single GPU.
+Since each cell of the patch is $30m\times 30m$, area of a $64\times 64$ patch is $1.423$ sq.
+miles.
+The entire eastern shore of Virginia has a total area of $2105$ sq miles, which would require $\sim 1480$ patches of size $64\times 64$.
+On a single GPU, making 36 hour forecasts for all these patches would take $2.88$ hours using Diff-Sparse.
+However, we can parallelize the forecasting of different patches.
+For instance, if we have $8$ GPUs available, making $36$ hour forecast for the entire Eastern Shore of Virginia would take $2.87/8\approx 0.36$ hours or $\sim 22$ minutes.
+Appendix E Inundation Scenario Queries
+Once Diff-Sparse is trained, we can use it to answer inundation scenario-based queries relevant to policymakers. In this Section, we discuss a few of such queries.
 Figure 6: (left) A school area in the Eastern Shore of Virginia.
@@
 A possible query is: what is the probability that the route coming out of the school will not be flooded in the next $T$ hours?
-Input: Query area $A$ as a Polygon, Flooding level threshold $d$, Prediction Length $T$, Trained Diff-Flood Model.
-1 $C_{A}\leftarrow$ Set of grid cells that overlap with $A$.
+Input: Query area AA as a Polygon, Flooding level threshold $d$, Prediction Length $T$, Trained Diff-Sparse Model.
+1 $C_{A}\leftarrow$ Set of grid cells that overlap with AA.
 2 $\mathcal{P}_{A}\leftarrow$ Set of patches where each patch $P\in\mathcal{P}_{A}$ contains at-least one cell $c\in C_{A}$.
@@
 […]
-return $p$.
+return pp.
 Algorithm 3 Calculate the probability that the flooding level in an area $A$ will be $\leq d$ units in the next $T$ hours.
@@
 Therefore, the probability that all cells $c\in C_{A}$ have flooding level $\leq d$ units is calculated to be $p=\prod_{P\in\mathcal{P}_{A}}probability(P_{\leq d})$ (line 3).
-    Since we require the probability of flooding level being above $d$ units in area $A$, our desired probability value is $(1-p)$.
+    Since we require the probability of flooding level being above $d$ units in area $A$, our desired probability value is (1p)(1-p).
   • Query 2: Given an evacuation route, what is the probability that the route will not be flooded in the next $T$ hours?
@@
 For instance, a more refined version of Query 1 can be: what is the probability that at-least half of the region $A$ will have flooding level above $d$ units in the next $T$ hours?
-The key idea for answering such queries is that, we have a sampling method in the form a trained Diff-Flood model and we can use it to sample inundation scenarios from the distribution of possible scenarios.
+The key idea for answering such queries is that, we have a sampling method in the form a trained Diff-Sparse model and we can use it to sample inundation scenarios from the distribution of possible scenarios.
 We use the sampled scenarios to calculate our desired probability value.
-Appendix D Hyper-parameters
+Appendix F Reproducibility Checklist Additional Information
+Our source code for Diff-Sparse and comparison with baseline methods are provided with the supplementary materials. The TideWatch inundation dataset will be provided upon request.
+Appendix G Hyper-parameters
 Hyper-parameter Value
 Learning rate 0.001
-Learning rate update step size 5 epochs
+Learning rate update policy Reduce on Plateau.
+Learning rate update patience 3 epochs.
 Learning rate update factor 0.5
-Number of epochs 30
+Number of epochs 40
 Training batch size 32
@@
 Diffusion $\beta_{max}$ 1
-NO. of scenarios generated for validation 2
-NO. of scenarios generated for test 8
+# of scenarios generated for validation 2
+# of scenarios generated for test 8
 Table 2: Shared hyper-parameters for all of our patch configurations in Table 1 of main paper.
@@
 Table 3: Hyper-parameters for different patch configurations
-Appendix E Patch Visualization
+Appendix H Patch Visualization
 Figure 7: Visualization of inundation on $10$, $64\times 64$ patches in the first $24$ hours of training data. The darker the shade of blue, the higher the inundation value.